### v1

# Managing Classical Processing Requirements for Quantum Error Correction

Abstract.
Quantum Error Correction requires decoders to process syndromes generated by the error-correction circuits. These decoders must process syndromes faster than they are being generated to prevent a backlog of undecoded syndromes that can exponentially increase the memory and time required to execute the program. This has resulted in the development of fast hardware decoders that accelerate decoding. Applications utilizing error-corrected quantum computers will require hundreds to thousands of logical qubits and provisioning a hardware decoder for every logical qubit can be very costly. In this work, we present a framework to reduce the number of hardware decoders and navigate the compute-memory trade-offs without sacrificing the performance or reliability of program execution. Through workload-centric characterizations, we propose efficient decoder scheduling policies which can reduce the number of hardware decoders required to run a program by up to 10$\times$ while consuming less than 100 MB of memory.
1. Introduction
As quantum computing enters a phase of rapid scaling to enable Fault-Tolerant Quantum Computing (FTQC), the classical processing resources required to support Quantum Error Correcting (QEC) codes must be scaled proportionally. QEC codes generate a stream of syndromes repeatedly by measuring parity qubits every cycle, and a decoder algorithm running on the classical control computer processes the stream of syndrome bits to detect errors and correct errors. Recent demonstrations have shown how the Surface Code (Fowler2012, ) can be deployed experimentally to suppress logical error rates (Google2023, ), how neutral atoms can be used to realize up to 48 logical qubits (Bluvstein2023, ), and how four logical qubits could be created with thirty physical qubits to achieve an 800$\times$ reduction in the error rate (quantinuum2024, ). These demonstrations are precursors to complex systems with more logical qubits requiring significant classical processing resources to enable fault-tolerant architectures.
[…]
Applications that can benefit from FTQC will require hundreds to thousands of logical qubits to function (Blunt2024, ). Depending on the error-correcting code used, replicating decoders for every logical qubit in the system can become very expensive and intractable in terms of cost and complexity. To reduce this cost, fast, hardware-efficient decoders have been proposed which sacrifice some accuracy for speed and scalability by making approximations in the decoding process (Ravi2023, ; Riverlane2023, ; Vittal2023, ; Alavisamani2024, ). However, catering to hundreds to thousands of logical qubits with these specialized decoders will still result in complex and costly systems – in this work, we aim to show how the total number of decoders can be reduced without affecting the performance or reliability of the quantum computer, thus allowing for more scalable classical processing for QEC.
Figure 1. (a) The compute and memory trade-offs for the classical processing required to implement QEC – fewer decoders result in an exponential increase in the memory required to store undecoded syndromes, and we aim to avoid this exponential growth in memory while reducing the number of decoders; (b) Virtual decoders – a scheduling policy will allocate decoders to the logical qubits.
[…]
[some figure]
In this paper, we present VQD: Virtual Quantum Decoding, a framework that aims to provide the illusion that there are decoders for every logical qubit while using significantly fewer hardware decoders to enable scalable and efficient classical processing necessary for fault-tolerant quantum computers. As shown in Fig. 1(a), the objective of VaDER is to reduce classical computing and memory resources needed to execute quantum programs on a fault-tolerant quantum computer without causing a performance slowdown or increase in the logical error rate. This is challenging because when we reduce the number of physical decoders, the memory required to store undecoded syndromes can grow exponentially if the syndrome generation and syndrome processing rates are not matched. More importantly, this exponential increase in memory due to syndrome backlog will lead to an exponential increase in the time it takes to process all the undecoded syndromes, thereby significantly increasing the execution time of the program (Terhal2015, ). Our experimental evaluations also show that if a logical qubit is not decoded for extended periods (which will occur if there are fewer decoders than qubits), then it can cause the decoder latency to increase due to an increase in undecoded errors. This increase in the decoder latency can affect the application of non-Clifford states – if decoding is delayed for a logical qubit before applying a non-Clifford state, the application of the non-Clifford state could be delayed since the decoder might take more time than usual to decode all prior rounds.
Given the challenges in sharing decoder hardware and to understand how it can be enabled, we characterized representative FTQC workloads to understand the decoding requirements from a performance and reliability perspective. Our characterization using a lattice surgery compiler (watkins2023high, ) revealed that there is a limited amount of operational parallelism due to long sequences of $T$ and $H$ gates resulting from Clifford + $T$ decomposition, which is necessary for universality. This is crucial, as non-Clifford gates are the reason real-time decoding is necessary for FTQC. Non-Clifford operations are the only operations where the decoding is in the critical path, and fortunately, they occur in a highly serialized manner. Therefore, a physical decoder per logical qubit is unnecessary and will lead to severe underutilization.
Armed with this insight, we propose a system architecture with significantly fewer physical decoders than the number of logical qubits. Furthermore, we design efficient decoder scheduling policies for such systems. Such a system can be visualized in Fig. 1(b). We propose a scheduling policy that minimizes the Longest Undecoded Sequence, termed as the MLS policy. We compare it with the Round Robin (RR) and Most Frequently Decoded (MFD) policies – our evaluations show that the MLS policy can reduce the number of hardware decoders by up to 10$\times$ while ensuring that no logical qubit remains undecoded for a significantly long period.
We also propose a noise-adaptive scheduling policy that can prioritize decoding of logical qubits that incur a sharp increase in the physical error rate due to phenomena such as cosmic rays (McEwen2021, ) and leakage due to heating (Miao2023, ). This involves a simple detector that can schedule decoding for a logical qubit in case the syndromes for that logical qubit show a sudden increase in bit-flips. Next, we show how some decoding tasks can be offloaded to software to further improve the efficacy of decoder scheduling policies.
Balancing compute and memory is a classic architectural problem, and we use VQD to explore these trade-offs. Prior research on decoders and classical processing required for fault-tolerant quantum computers have focused on reducing the hardware cost of implementing decoders by making approximations in the decoding algorithm, sometimes at the cost of accuracy (Vittal2023, ; Riverlane2023, ; Alavisamani2024, ). With VaDER, we show that even if individual decoders have a high hardware cost, the overall cost can be reduced significantly by virtualizing decoders.
2. Quantum Error Correction and Decoding
In this section, we cover high-level details of Quantum Error Correction and the role of decoders.
2.1. Quantum Error Correction
Quantum Error Correction (QEC) improves the reliability of a system by utilizing many physical qubits to encode a single logical qubit (Shor1995, ; Knill1997, ). Most QEC codes can be categorized as stabilizer codes (Terhal2015, ) – some promising stabilizer codes include quantum Low Distance Parity Check (qLDPC) codes (Bravyi2024, ) and Surface Codes (Fowler2012, ). Owing to their relatively relaxed connectivity requirements that can be realized with hardware available today, we focus specifically on the Surface Code. Note that this work can be extended to other QEC codes apart from the Surface Code.
Figure 2. (a) A logical qubit ($d=3$); (b) Syndrome generation and measurements; (c) A typical procedure for detecting errors by decoding syndromes.
[…]
[some figure]
A single rotated Surface Code patch of distance $d=3$ is shown in Fig. 2(a) (Horsman2012, ). Black circles denote data qubits and each data qubit is connected to $X$ and $Z$ parity qubits. The Surface Code works by repeatedly measuring syndromes, which correspond to the measurements performed on all $X$ and $Z$ measure qubits in a patch after performing the sequence of gates shown in Fig. 2(b). By repeatedly measuring these syndromes, both bit-flip and phase-flip errors occurring within a patch can be detected by the decoder.
2.2. Decoding Errors
Fig. 2(c) shows the general procedure of how any generic QEC code works – syndromes are constantly being generated, which are then fed to a decoder. Syndromes contain information about which qubits have flipped in every round of syndrome measurements, and these flips allow the decoder to determine what errors on the data qubits caused those flips. Since errors can always be expressed in the form of Pauli gate, they can be corrected in software without executing any physical operations on the logical qubits. This is achieved by updating the Pauli frames for all data qubits that make up that logical qubit, which adjusts the interpretation of future measurements by accounting for the error that was detected (Terhal2015, ). For the Surface Code, the decoding problem is commonly formulated as a Minimum Weight Perfect Matching (MWPM) problem, which leverages a graph representation of the syndrome measurements (Fowler2015, ; Wu2022, ).
Figure 3. Consumption of a magic ($T$) state with LS.
[…]
[some figure]
2.3. Non-Clifford Gates
On a Surface Code error-corrected quantum computer, all Clifford gates can be performed reliably either in software or via logical operations performed via Braiding (Fowler2012, ) or Lattice Surgery (Horsman2012, ). However, non-Clifford gates such as the $T$ gate cannot be applied in a fault-tolerant manner directly. This is because logical qubits initialized with the $T$ gate will have an error probability equal to the underlying physical error rate of the system, $p$, thus making them impure (Litinski2019magic, ). However, multiple impure states can be used to distill fewer, purer logical qubits with a $T$ state (known as a magic state $|m\rangle$) – this process is known as magic state distillation (Bravyi2012, ; Gupta2024, ).
Fig. 3 shows how a $T$ gate can be applied to a logical qubit $P$ in a fault-tolerant manner by using Lattice Surgery (LS) (Litinski2019, ). The magic state $|m\rangle$ is a purified $T$-state. Since a non-Clifford gate is being applied, all prior errors that affected $P$ must be known before $|m\rangle$ is applied to prevent errors from spreading (Terhal2015, ). Lattice Surgery can be used to perform a $Z\otimes Z$ operation on $P$ and $|m\rangle$ to apply the magic state (Litinski2019, ; Litinski2019magic, ). Once the $Z\otimes Z$ operation is performed, the decoding result of $P$ prior to Lattice Surgery is combined with the decoding result of Lattice Surgery multi-body measurement and the measurement of the patch containing the magic state $|m\rangle$ to determine an appropriate Clifford correction [^1] The auto-corrected $\pi/8$ gate in (Litinski2019, ; Litinski2019magic, ) uses an additional ancillary qubit that has not been shown.. This correction needs to be known before the next logical operation involving a non-Clifford gate.
[…]
[some figure]
Figure 4. (a) Exponential increase in the memory required to store undecoded syndromes for different number of available decoders for the wstate-60 benchmark – 200 decoders correspond to assigning a decoder to every logical qubit; (b) Number of $T$-gates required for different workloads; (c) Exponential increase in the decoder latency per round (in nanoseconds) as the number of rounds is increased from $d$ to $20d$ – the increase is higher for larger code distances ($p=10^{-4}$); (d) Slow increase in the logical error rate as the number of rounds of error correction increase with $p=10^{-3}$, $p=10^{-4}$; (e) Histogram of the number of concurrent critical decodes for different workloads with the EDPC layout – there is limited parallelism as far as critical decodes are concerned (f) Number of FPGAs required for decoders to implement different workloads with (i) a Single $T$-factory and (ii) Optimal number of $T$-factories.
[…]
[fig]
2.4. Critical Decodes
For a logical qubit that is only executing Clifford gates, errors can be decoded at any point of time (even after the experiment has ended). This is because all Clifford corrections can be commuted to the end of the circuit, essentially allowing the syndromes to be post-processed rather than decoded in real-time (Terhal2015, ). However, universal fault-tolerant quantum computer require non-Clifford gates such as the $T$-gate – this makes real-time decoding a necessity since syndromes for a patch must be decoded before the application of a non-Clifford gate. We call decodes that must happen before the application of a non-Clifford gate critical decodes since all syndromes generated up to that point for that logical qubit must be decoded before computation can proceed.
3. Classical Processing Requirements
Having explained the relevant details about QEC and the role of decoders and critical decodes, we now cover the classical processing requirements for FTQC.
3.1. Syndrome Generation and Processing Rates
As discussed in Section 2, applying non-Clifford gates requires decoders to be up-to date with the latest syndrome for a logical qubit before computation can proceed. As shown by Terhal (Terhal2015, , p. 20), if the rate at which syndromes are generated $r_{gen}$ is faster than the rate at which they are processed $r_{proc}$, the memory required for storing syndromes that are yet to be decoded increases exponentially (referred to as the backlog problem). This exponential increase in the memory required also leads to an exponential increase in the runtime of the workload. Fig. 4(a) shows the exponential increase in the memory required to store undecoded syndromes for the wstate-60 benchmark with the number of rounds for various numbers of available decoders. For larger and longer running workloads, the memory requirements will be much higher. By reducing the number of available decoders, $r_{proc}$ is effectively reduced, leading to the exponential growth in memory. Decoding is necessary at runtime when consuming $T$-states, as discussed in Section 2, and Fig. 4(b) shows that it will be a frequent operation for most workloads.
$r_{proc}$ must be consistently greater than $r_{gen}$ to prevent an exponential increase in memory requirements.
3.2. Why is Real-Time Decoding Needed?
[…]
Leaving syndromes undecoded for even tens of rounds can result in an exponential slowdown in the decoder processing rate.
3.3. Concurrency and Delayed Decoding
Critical decodes must be serviced during the execution of a program to avoid the increase in memory requirements and processing times discussed above. If there are many concurrent critical decodes that occur frequently during the execution of a program, an appropriate number of decoders will be needed to process these critical decodes. The level of concurrency depends entirely on the layout used to build the quantum computer – layouts such as the Compact (Litinski2019, , p. 7), Fast (Litinski2019, , p. 9), and the Edge-Disjoint Path (EDPC) (Beverland2022edpc, ) are some proposed layouts that can be used to build fault-tolerant quantum computers using the Surface Code. Layouts determine the number of physical qubits required – for example, the Compact layout requires the fewest physical qubits since there is a single routing lane at the cost of completely serializing operations. The Fast and EDPC layouts allow more concurrency at the cost of more physical qubits.
What is the average level of concurrency when executing a quantum program on a fault-tolerant quantum computer? Fig. 4(e) shows a histogram of the number of critical decodes [^3] For the Surface Code, we assume every logical qubit requires two decoders – one each for the $X$ and $Z$ observables. for select workloads generated by the Lattice Surgery Compiler (watkins2023high, ). This histogram shows that the peak concurrency is attained very infrequently. This implies that most logical qubits function as memory qubits or execute Clifford gates more often than $T$-gates, and this can allow some qubits to not be decoded in real-time.
Quantum programs are serial in terms of $T$-gates applied – not every qubit always requires access to a fast decoder.
3.4. Goal: Make Classical Processing Efficient
Provisioning a hardware decoder for every logical qubit in the system can be resource intensive – Fig. 4(f) shows an estimate of the number of FPGAs required just for decoding when (i) a single distillation factory is used, and (ii) an optimal number of distillation factories are used for different workloads ($\sim$10FPGA LUTs/decoder (Vittal2023, ; Riverlane2023, )). This estimation was done using the Azure QRE (Beverland2022, ) that uses the Fast layout. Note that the total hardware requirement will be significantly higher because of control and readout components. Having shown how the syndrome processing rate is crucial in ensuring that computation does not require excessive memory and time and how quantum programs are inherently serial in terms of critical decodes, the question we seek to answer in this work is–
How can we minimize the use of hardware decoders and lower classical processing costs without sacrificing the performance and reliability?
4. Virtual Quantum Decoders
We now show how the number of hardware decoders can be reduced to be less than the number of logical qubits in the system, and how decoding can be scheduled to prevent excessive accumulation of undecoded syndromes.
Figure 5. (Left) Decoders for every logical qubit; (Right) Time-division multiplexing of decoders between qubits.
[…]
[some figure]
4.1. Working with Fewer Hardware Decoders
Reducing the number of available decoders implies that qubits will share hardware resources, resulting in time-division multiplexing of decoder instances among logical qubits. Fig. 5 shows how compared to a system with decoders for every logical qubit in the system ($N$ qubits, $N$ decoders), a system with fewer ($M$, $N>M$) decoders will require resources to be shared with time.
Time-division multiplexing of hardware resources will require the following considerations: (i) If the number of critical decodes at a given time step exceed the number of hardware decoders, the overflowing critical decodes will have to be deferred to the next available time step, and (ii) Qubits cannot be left undecoded for extended periods of time. For the first consideration, deferring critical decodes will increase serialization in the program – this offsets all benefits offered by the Fast and EDPC layouts. For the second consideration, leaving a qubit undecoded for too long will result in an exponential increase in the syndrome processing latency and memory required to store undecoded syndromes.
Since not all qubits will be involved in critical decodes at every time step, there will be some decoders available at a given point of time which will not be decoding a logical qubit involved in the consumption of a $T$-state. Allocating these free hardware decoders to logical qubits at every time thus becomes a scheduling problem.
Before we discuss different decoder scheduling policies, it is important to understand the time granularity at which any scheduling policy will operate on. Since logical operations (Clifford or non-Clifford) in a Surface Code error-corrected quantum computer will require at least $d$ rounds before the next operations, we define a slice (watkins2023high, ) as the smallest time step between logical operations that a decoder scheduling policy can work on. Every slice consists of $d$ rounds of syndrome measurements, thus making the scheduling policy agnostic of the actual code distance used.
Figure 6. (a) Illustration of the longest undecoded sequence – $Q_{1}$ has the longest undecoded sequence before the last decode; (b) MFD policy – undecoded qubits are sorted according to the number of critical decodes they are involved in after time slice $t$ and $M-C$ qubits are selected from this sorted list; (c) RR policy – the $M-C$ qubits decoded in time slice $t$ are not decoded in time slice $t+1$; (d) MLS policy – undecoded qubits are sorted according to their undecoded sequence lengths and $M-C$ qubits are selected from this sorted list.
[…]
[some figure]
4.2. Decoder Scheduling Policies
Static decoder scheduling refers to decoder scheduling that can be performed at compile time. Since most quantum programs do not have any control-flow instructions, scheduling can be performed statically. Scheduling decoders is similar to CPU scheduling performed by all operating systems today, where the number of processes is more than the number of available processor cores (Liu1973, ).
Longest Undecoded Sequence: To quantify the fairness of a decoder scheduling policy, we use ‘Longest Undecoded Sequence’, which measures how well the decoders are servicing all logical qubits. A large undecoded sequence length implies that a qubit has been left undecoded for a long time – increasing the memory consumed to store undecoded syndromes. Fig. 6(a) shows an example of determining the longest undecoded sequence length.
Consider an arbitrary time slice $t$ in the execution of a quantum program. There are $N$ logical qubits and $M$ hardware decoders ($N>M$). All decoding scheduling policies will have two components: The first will assign the decoders necessary for all critical decodes $C$ in the time slice $t$. The second will assign all the remaining $M-C$ hardware decoders to the $N-C$ qubits based on the scheduling policy used. We now discuss three decoder scheduling policies (all policies are illustrated in Fig. 6(b) – Fig. 6(d)):
4.2.1. Most Frequently Critically Decoded (MFD)
A logical qubit that consumes a significant number of $T$-states during the execution of a program would have a frequent requirement of critical decodes – leaving such a logical qubit undecoded for more than a few slices would make subsequent critical decodes take longer, thus slowing down computation. This motivates the MFD scheduling policy that prioritizes decoding of logical qubits that have numerous critical decodes in the future at any given time slice. The MFD policy will ensure that future critical decodes have a minimized number of undecoded syndromes for the qubits that have frequent critical decodes.
Caveats: Because the MFD policy prioritizes logical qubits with frequent critical decodes, it will likely starve other qubits of decoding, leading to longer undecoded sequences.
4.2.2. Round Robin (RR)
Derived from CPU scheduling policies used by operating systems, the RR policy does not prioritize any specific logical qubits – rather, it chooses $M-C$ qubits in a round-robin manner in every time slice to ensure fairness for all qubits in the system.
Caveats: For regions of a program where there are many critical decodes, the RR policy could still starve some logical qubits since $M-C$ will be much smaller, yielding a smaller window for decoders to be assigned. Since there is no prioritization, the RR policy will not be able to rectify this until the round-robin window reaches the qubits being starved.
4.2.3. Minimize Longest Undecoded Sequence (MLS)
The longest undecoded sequence length at any given time slice is an indicator of how well the decoder scheduling policy is servicing all qubits in the system. We use this as a motivator for the MLS policy, which tries to minimize the longest undecoded sequence at every time slice. The MLS policy works as follows: at any time slice $t$, qubits are sorted on the basis of their current undecoded sequence lengths. Then, $M-C$ qubits with the largest undecoded sequence lengths are assigned hardware decoders.
Caveats: In cases where there the number of logical qubits is far greater than the number of decoders ($N>>M$), the MLS policy will not be able to work effectively.
4.3. Noise-Adaptive Decoder Scheduling
While control-flow instructions would necessitate runtime scheduling of decoders, events such as cosmic rays (McEwen2021, ) and leakage (Miao2023, ) can result in a temporary burst of errors for some physical qubits in the lattice that can impact some logical qubits. Scheduling after a control-flow instruction can be performed using any static scheduling policies for the program after the control-flow instruction. However, since the static scheduling policies do not account for the error-rate, spikes in errors due to cosmic rays and leakage cannot be factored at runtime without hardware support.
Figure 7. (a) Increase in the average number of bit-flips in syndromes for different code distances as the physical error rate is increased; (b) An error burst results in increased bit-flips in the syndromes of patch $P$ which can be detected and used to prioritize the decoding of $P$ in the next time step; (c) Decoding can be offloaded to software provided there is enough time before a critical decode occurs.
[…]
[fig]
Detecting a spike in the physical error rate will either require errors to be decoded or additional hardware modules to detect the spike. While error-correcting codes can tolerate temporary increases in the error-rate (Google2023, ), the increase in errors can result in longer decoding latencies since the decoding task becomes harder with more errors. If a logical qubit affected by these events is not scheduled for decoding immediately after the event, decoding it before applying a non-Clifford gate could take longer, thus delaying the operation and causing a slowdown. As shown in Fig. 7(a), an increase in the physical error rate results in a higher number of bit-flips (especially for larger code distances), which can be detected with simple components in the control hardware. Fig. 7(b) shows how additional flips can be detected and used to dynamically prioritize the decoding of an arbitrary patch $P$, which suffers from a temporary burst of errors. Note that the detection is different from decoding – we are merely predicting that there are more errors due to higher bit-flips in the syndromes.
4.4. Offloading to Software Decoders
Software decoders are slow and also have a higher variance in decoding latencies (Delfosse2023, ). However, when scheduling decoding tasks for logical qubits, software decoders can be leveraged to further reduce the undecoded sequence lengths. As shown in Fig. 7(c), some syndromes for a logical qubit can be offloaded to software while the hardware decoders are busy elsewhere. To prevent scheduled hardware decoding from being delayed, a buffer (three slices in this example) must be used to ensure that the software offloading completes before the next hardware decode.
5. Decoding for Distillation Factories
The decoder scheduling policies in the previous sections catered only to the decoding of algorithmic logical qubits (data logical qubits, magic state storage, ancillary logical qubits required for Lattice Surgery). In this section, we discuss decoding for distillation factories.
5.1. Distillation Factories
Magic state distillation factories generate few low-error logical qubits with non-Clifford states from many high-error logical qubits. Distillation factories run for very short periods at a time – this allows for smaller code distances to be used for creating the logical qubits for distillation (Litinski2019magic, ). As shown in Fig. 8, the error probability of a magic state is low enough to be useful even with $d=7$ ($d$ refers to $d_{X}$ in (Litinski2019magic, )).
Figure 8. Output error probabilities of magic states after 15-1 distillation for different $d$ (Litinski2019magic, ) ($p=10^{-4}$).
[…]
[fig]
5.2. Using Fast, Low-Footprint Decoders
Smaller code distances for distillation factories provide two main benefits: the number of physical qubits required are much lower, and more importantly, both hardware (Vittal2023, ; Smith2023, ; Delfosse2020, ; Caune2023, ) and software (Delfosse2023, ) decoders are faster and less complex. For example, LUT based decoders (Das2022lilliput, ) have been shown to be effective up to $d=5$ without requiring significant hardware resources. Predecoders (Delfosse2020, ; Ravi2023, ; Smith2023, ) reduce the complexity and decoding effort required for lower code distances as well. Compared to algorithmic logical qubits, which require a large code distance to survive millions of error correction cycles (Beverland2022, ; Blunt2024, ), the decoding requirements of distillation factories are far more relaxed, which reduces the hardware resource requirements as well.
The decoding overhead of magic state distillation is significantly lower than algorithmic logical qubits – it can thus leverage lightweight decoders, considerably reducing the hardware cost for distillation factories.
6. Methodology
We now describe the methodology used to evaluate different decoder scheduling policies and for estimating classical resources required for executing workloads on a Surface Code error-corrected quantum computer.
6.1. Compiler
We use the Lattice Surgery Compiler (LSC) (watkins2023high, ) to generate Intermediate Representations (IR) of workloads that can be executed on an error-corrected quantum computer using the Surface Code with lattice Surgery. LSC can generate IR that denote Lattice Surgery instructions from the QASM (qasm, ) representation of a workload. LSC handles mapping and routing based on the layout provided to the compiler. We configure LSC to use a ‘wave’ scheduling that maximizes the number of concurrent instructions executed in every time slice. LSC also uses Gridsynth (gridsynth, ) to deal with arbitrary rotations. However, since it is still under development, LSC has some limitations:
  • LSC is limited to multi-body measurements between only two logical qubits.
  • LSC abstracts away distillation factories, only magic state storage sites are considered.
  • LSC works for a limited set of layouts and is extremely slow for large workloads like shor.
6.2. Simulation Framework
Using the IR generated by LSC, we build a framework that can parse the IR and determine the critical decodes in every slice, generate a timeline of all operations, and assign decoders to all logical qubits depending on the scheduling policy. In case the number of critical decodes in a particular slice are more than the number of hardware decoders configured, decoder-resources can rewrite the IR to defer critical decodes to the next slice (potentially increasing the execution time of the program).
Layouts: We use three layouts for our evaluations – Fast and Compact layouts (Litinski2019, ), and the EDPC layout (Beverland2022edpc, ). The Compact layout uses the fewest logical qubits and, due to a single routing lane, allows only one magic state to be consumed per time slice – we thus use it only to compare total execution times with the Fast and EDPC layouts.
Benchmarks: We use benchmarks from MQT Bench (quetschlich2023mqtbench, ) and QASMBench (Li2023, ). We use shor-15, a chemistry workload gndstate-14, a NISQ workload qaoa-14444Used for its arbitrary rotations and similarity to chemistry workloads., random, wstate, and arithmetic workloads adder-28, multiplier-45, Quantum Fourier Transform qft-20 – which can be used as building blocks for other algorithms.
Other Software: Stim (Gidney2021, ) was used for simulating stabilizer circuits to generate syndromes and error rates. Azure QRE (Beverland2022, ) was used for resource estimations.
7. Evaluations
In this section, we present some results for different scheduling policies and savings in decoder hardware.
Figure 9. (a) Max. concurrent critical decodes for all layouts and workloads; (b) Total execution time (in code cycles) needed to execute all workloads with different layouts (c) Estimated number of logical qubits for all workloads (code distances estimated by Azure QRE (Beverland2022, ) also annotated).
[…]
[some figure]
7.1. Research Questions
We aim to answer the following questions:
  1. (1)
    How many decoders can we virtualize in the system without affecting performance?
  2. (2)
    How long do qubits go undecoded when using different scheduling policies?
  3. (3)
    How do scheduling policies affect memory usage when using virtualized decoders?
7.2. Baseline Statistics
[…]
7.3. Decoder Scheduling Efficacy
Figure 10. Hardware decoders used for different configurations (EDPC layout).
[…]
[some figure]
(a) Max. Concurrency
(b) Midpoint
Figure 11. Longest undecoded sequence when using the (a) Max. Concurrency; (b) Midpoint configurations.
[…]
[some figure]
Fig. 10 shows the three decoder configurations selected for this work.
  • The All Qubits configuration denotes the baseline where all qubits have a decoder.
  • Max. Concurrency refers to the configuration where the number of hardware decoders in the system corresponds to the peak concurrent critical decodes for every workload shown in Fig. 9(a).
  • Midpoint refers to a configuration where the number of hardware decoders is the midpoint between the max. and min. concurrent critical decodes ($=\frac{Max.+Min.}{2}$).
The minimum concurrent critical decodes corresponds to two critical decodes between two logical qubits. All evaluations are for algorithmic logical qubits, logical qubits required for distillation are not considered. For the All Qubits configuration, the longest undecoded sequence length will be zero, since every logical qubit has an assigned decoder.
7.3.1. Longest Undecoded Sequences
To evaluate the performance of the decoder scheduling policies described in Section 4, we determine the longest undecoded sequence lengths for all workloads when using the Max. Concurrency and Midpoint configurations. Since these configurations use far fewer hardware decoders than qubits, the longest undecoded sequence is a good measure of whether qubits are being starved of decoding. Fig. 11(b) shows the longest undecoded sequence lengths for the (a) Max. Concurrency and (b) Midpoint configurations. The MFD policy leads to qubits being starved of decoding, since it prioritizes qubits that have frequent critical decodes. For the Max. Concurrency configuration, the gndstate and qaoa workloads do relatively well with the MFD configuration signifying that almost all logical qubits have a similar number of critical decodes, leading to a fairer scheduling. While the RR policy performs significantly better than the MFD policy, the MLS policy consistently performs better than both policies – MLS reduces the longest undecoded sequence lengths for almost every workload to $\sim 10$ slices for both Max. Concurrency and Midpoint configurations.
(a) Max. Concurrency
(b) Midpoint
Figure 12. Peak memory required for storing undecoded syndromes for different scheduling policies when using the (a) Max. Concurrency; (b) Midpoint configurations.
[…]
[some figure]
7.3.2. Memory Usage
The reduction in the longest undecoded sequence lengths also corresponds to lower memory usage for storing undecoded syndromes. This is crucial since reducing the number of hardware decoders will require more memory to store syndromes for qubits that have not been decoded. Fig. 12(b) shows the memory required for different workloads with the (a) Max. Concurrency and (b) Midpoint configurations (the Azure QRE estimated the code distances used to determine the memory requirements). Due to longer undecoded sequences, the MFD policy can require up to 100 GB of memory for some workloads while the MLS policy rarely requires more than 100 MB of memory, which is orders of magnitude better than the MFD policy and 2-4x better than the RR policy.
Figure 13. Total slices required by workloads with different decoder configurations normalized to the slices required by the baseline (MLS policy).
[…]
[some figure]
7.3.3. Slowdown due to Fewer Decoders
Fewer hardware decoders imply that some critical decodes in an arbitrary slice have to be deferred to subsequent slices, resulting in potentially more slices for completing the program. Fig. 13 shows the number of slices required for all workloads normalized with respect to the baseline number of slices shown in Fig. 9(b) for the Min. Concurrency, Max. Concurrency, and Midpoint configurations. Since the Min. Concurrency configuration allocates only four decoders for two critical decodes per slice, some workloads are slowed down by ¿10%. The Midpoint configuration however does not cause any slowdown except in the gndstate workload, and there is no slowdown for the Max. Concurrency configuration.
Figure 14. Increase in the LER (lower is better) with the Midpoint configuration.
[…]
[some figure]
7.3.4. Impact on Logical Error Rate
As discussed in Section 3, delaying decoding does not affect the logical error rate (LER) by itself. However, leaving a qubit undecoded for a long time before a critical decode can slow the decoder down, thus requiring additional rounds to complete the computation. In this evaluation, we would thus like to show how the longest undecoded sequence length can impact the LER. Fig. 14 shows how the MFD policy can increase the final LER of the computation due to longer undecoded sequence lengths. The RR policy increases the LER slightly for adder-28 and random-40, and the MLS policy does not incur any degradation in the LER. Note that this estimation is optimistic since we assumed the number of rounds increases linearly with the undecoded sequence length – in reality, it could be worse since the decoder latency can increase exponentially with the undecoded sequence length.
Figure 15. Reduction in the peak memory usage when using the Midpoint configuration with software offloading.
[…]
[some figure]
7.3.5. Software Offloading
The longest undecoded sequences (and consequently the memory requirements) can be reduced further by leveraging software decoders. The only constraint while doing so is that due to longer software decoding latencies, critical decodes should not be delayed because prior software decodes for a logical qubit have not yet finished. For evaluating the effect of software decoding, we set the number of hardware decoders to the Midpoint configuration and make a pessimistic assumption that a single slice worth of syndromes takes three slices (about 3$\times$×d microseconds) to be decoded in software (in reality, it could be much lower with optimized software decoders (Higott2023, )). Fig. 15 shows the reduction in the peak memory usage for all scheduling polices when software offloading is performed – software offloading can achieve a reduction of up to 3x.
7.4. Discussion
The results shown in previous sections show that VQD reduces the number of hardware decoders for algorithmic logical qubits by nearly one order of magnitude for most workloads with the Midpoint configuration, which, when combined with the MLS scheduling policy, results in significantly reduced memory requirements and low undecoded sequence lengths. Memory requirements and undecoded sequence lengths can be further reduced by offloading some non-critical decodes to software.
[…]
Applicability to other codes: While our evaluations have focused on the Surface Code, other error-correcting codes such as quantum LDPC codes (Bravyi2024, ) also require decoding with a significantly higher complexity than Surface Code. qLDPC codes use Belief Propagation for decoding errors (Old2023, ) along with Ordered Statistics Decoding (OSD) (Roffe2020, ; Fuentes2021, ) to enable accurate decoding. These algorithms are complex and highly resource intensive (Gong2024, ). Building fast, accurate decoders for such codes will likely require significant hardware resources. Our work enables amortizing the cost of expensive decoders via virtualization to build efficient and scalable quantum memory using qLDPC codes.
Better Capacity Planning: We envision large-scale quantum computers will be closely integrated with HPC-style systems, where scientific applications can leverage quantum subroutines using QPUs (humbleQPU, ). In this setting, non-critical software decoders can run on traditional HPC platforms to alleviate the pressure on hardware decoders. Moreover, the virtualization of decoders can help us harness shot-level parallelism – all quantum programs, even on FTQC, must be executed multiple times. We can concurrently run the copies of quantum programs on multiple QPUs. However, quantum resources increase linearly for running “$k$” copies concurrently. Our work, VQD, shows that with decoder virtualization, we can enable effective sharing of classical resources, dramatically reducing overall costs and improving resource utilization.
8. Related Work
This is one of the first works to perform a workload-oriented study of the classical processing requirements and system-level scheduling policies for error-corrected quantum computers. Prior to this work, Bombín et al. (Bombin2023, ) introduced modular decoding, which is the closest work that divides the global decoding task to sub-tasks without sacrificing decoder accuracy. However, this work, and other works such as parallelized window decoding (Skoric2023, ; Tan2022, ) always assume decoders for every logical qubit. This work shows that not all qubits require access to fast decoders at all times, thus allowing decoders to be virtualized. Other works that are broadly connected to this work are summarized below.
System-level Studies: Delfosse et al. (Delfosse2023, ) studied the speed vs. accuracy tradeoff for decoders used in FTQC. XQSim (Byun2022, ) is a full-system FTQC simulator. Stein et al. (Stein2023, ) proposed a heterogeneous architecture for FTQC, virtual logical qubits were proposed in (Baker2021, ), Lin et al. (Lin2023, ) explored modular architectures for error-correcting codes and scheduling for distillation factories was proposed in (Ding2018, ). (Kim2024, ) described a blueprint of a fault-tolerant quantum computer.
Decoder Designs: Neural network based decoders (googleRNN, ; ueno2022, ; Overwater2022, ; Meinerz2022, ; Gicev2023, ; Varsamopoulos2017, ), LUT-based decoders (Das2022lilliput, ; Tomita2014, ), decoders based on the union-find algorithm (Das2022afs, ; Riverlane2023, ), and optimized MWPM decoders (Vittal2023, ; Alavisamani2024, ) have been proposed. In general, neural network decoders are generally far slower and therefore not ideal for fast qubit technologies such as superconducting qubits. Other predecoders (Delfosse2020, ; Smith2023, ) and partial decoders (Caune2023, ) have also been proposed. Decoders based on superconducting logic (Ueno2021, ; Ravi2023, ) target cryogenic implementations.
9. Conclusions
Scaling quantum computers to enable Quantum Error Correction will require specialized hardware for decoding errors. Prior work has focused on reducing the hardware resources required to build decoders. In this work, we take a full-system view and show that with the right decoder scheduling policy, it is not necessary for an error-corrected quantum computer to provide every logical qubit with a dedicated hardware decoder. The MLS policy enables the reduction of hardware decoders by up to 10x while requiring $\sim$100 MB or less of memory for storing undecoded syndromes without increasing the program execution time or the target logical error rate. The efficacy of the MLS policy is enhanced with software offloading of some decoding tasks. We also propose a noise-adaptive scheduling mechanism that can prioritize the decoding of logical qubits that incur a temporary increase in the physical error rate.

### v2

@@
 Quantum Error Correction requires decoders to process syndromes generated by the error-correction circuits.
-These decoders must process syndromes faster than they are being generated to prevent a backlog of undecoded syndromes that can exponentially increase the memory and time required to execute the program.
-This has resulted in the development of fast hardware decoders that accelerate decoding.
+These decoders must process syndromes faster than they are being generated to prevent a backlog of undecoded syndromes.
+This backlog can exponentially increase the time required to execute the program, which has resulted in the development of fast hardware decoders that accelerate decoding.
 Applications utilizing error-corrected quantum computers will require hundreds to thousands of logical qubits and provisioning a hardware decoder for every logical qubit can be very costly.
-In this work, we present a framework to reduce the number of hardware decoders and navigate the compute-memory trade-offs without sacrificing the performance or reliability of program execution.
-Through workload-centric characterizations, we propose efficient decoder scheduling policies which can reduce the number of hardware decoders required to run a program by up to 10$\times$ while consuming less than 100 MB of memory.
+In this work, we present a framework to reduce the number of hardware decoders and navigate the compute-performace trade-offs without sacrificing the performance or reliability of program execution.
+Through workload-centric characterizations performed by our framework, we propose efficient decoder scheduling policies that can reduce the number of hardware decoders required to run a program by up to 10$\times$.
+With the proposed framework, scalability can be achieved via decoder virtualization, and individual decoders can be built to maximize accuracy and performance.
+Virtualization, Quantum Error Correction, Decoders
+conference: The 52nd IEEE/ACM International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japancopyright: none
 1. Introduction
 As quantum computing enters a phase of rapid scaling to enable Fault-Tolerant Quantum Computing (FTQC), the classical processing resources required to support Quantum Error Correcting (QEC) codes must be scaled proportionally.
-QEC codes generate a stream of syndromes repeatedly by measuring parity qubits every cycle, and a decoder algorithm running on the classical control computer processes the stream of syndrome bits to detect errors and correct errors.
-Recent demonstrations have shown how the Surface Code (Fowler2012, ) can be deployed experimentally to suppress logical error rates (Google2023, ), how neutral atoms can be used to realize up to 48 logical qubits (Bluvstein2023, ), and how four logical qubits could be created with thirty physical qubits to achieve an 800$\times$ reduction in the error rate (quantinuum2024, ).
-These demonstrations are precursors to complex systems with more logical qubits requiring significant classical processing resources to enable fault-tolerant architectures.
+QEC codes generate a stream of syndromes repeatedly by measuring parity qubits every cycle, and a decoder algorithm running on the classical control computer processes the stream of syndrome bits to detect and correct errors.
+Recent demonstrations have shown the Surface Code (Fowler et al., 2012) can be deployed experimentally to suppress logical error rates (Acharya et al., 2024, 2023), creating 48 logical qubits using neutral atoms (Bluvstein et al., 2023), an 800$\times$ reduction in the error rate using four logical qubits (da Silva et al., 2024).
+These demonstrations are precursors to complex systems with more logical qubits that will require significant classical resources to enable FTQC systems.
 […]
-Applications that can benefit from FTQC will require hundreds to thousands of logical qubits to function (Blunt2024, ).
+Applications that can benefit from FTQC will require hundreds to thousands of logical qubits to function (Blunt et al., 2024).
 Depending on the error-correcting code used, replicating decoders for every logical qubit in the system can become very expensive and intractable in terms of cost and complexity.
-To reduce this cost, fast, hardware-efficient decoders have been proposed which sacrifice some accuracy for speed and scalability by making approximations in the decoding process (Ravi2023, ; Riverlane2023, ; Vittal2023, ; Alavisamani2024, ).
+This is especially because decoders add to the classical hardware already needed for qubit control and readout.
+To reduce this cost, fast, hardware-efficient decoders have been proposed which sacrifice some accuracy for speed and scalability (Ravi et al., 2023; Barber et al., 2023; Vittal et al., 2023; Alavisamani et al., 2024).
 However, catering to hundreds to thousands of logical qubits with these specialized decoders will still result in complex and costly systems – in this work, we aim to show how the total number of decoders can be reduced without affecting the performance or reliability of the quantum computer, thus allowing for more scalable classical processing for QEC.
 Figure 1.
-(a) The compute and memory trade-offs for the classical processing required to implement QEC – fewer decoders result in an exponential increase in the memory required to store undecoded syndromes, and we aim to avoid this exponential growth in memory while reducing the number of decoders; (b) Virtual decoders – a scheduling policy will allocate decoders to the logical qubits.
+(a) Fewer decoders result in a slowdown in the computation due to large backlogs of undecoded syndromes.
+VQD avoids this slowdown in computation while reducing the number of decoders; (b) Multi-objective optimization problem for designing decoders without any virtualization – current decoders are designed to minimize latency and maximize scalability, accuracy; (c) Virtualization allows the optimization problem to be reduced to just speed and accuracy – decoder virtualization can improve scalability.
 […]
 [some figure]
-In this paper, we present VQD: Virtual Quantum Decoding, a framework that aims to provide the illusion that there are decoders for every logical qubit while using significantly fewer hardware decoders to enable scalable and efficient classical processing necessary for fault-tolerant quantum computers.
-As shown in Fig.
-1(a), the objective of VaDER is to reduce classical computing and memory resources needed to execute quantum programs on a fault-tolerant quantum computer without causing a performance slowdown or increase in the logical error rate.
-This is challenging because when we reduce the number of physical decoders, the memory required to store undecoded syndromes can grow exponentially if the syndrome generation and syndrome processing rates are not matched.
-More importantly, this exponential increase in memory due to syndrome backlog will lead to an exponential increase in the time it takes to process all the undecoded syndromes, thereby significantly increasing the execution time of the program (Terhal2015, ).
+In this paper, we present VQD: Virtual Quantum and Decoding, a framework that aims to provide the illusion that there are decoders for every logical qubit while using significantly fewer hardware decoders to enable scalable and efficient classical processing necessary for fault-tolerant quantum computers.
+Reducing the number of decoders is not straightforward – if there are fewer decoders than logical qubits, some logical qubits will remain undecoded for some periods of time.
+When a logical qubit remains undecoded for some rounds of syndrome generation until a non-Clifford gate is executed on it, the overall decoding task is more than the original set of undecoded syndromes since syndromes keep getting generated as the decoder processes the original set of undecoded syndromes.
+This introduces a slowdown in the computation, as shown in Figure 1(a).
+The goal of this work is to avoid the slowdown in computation while reducing the compute resources required for decoding.
 Our experimental evaluations also show that if a logical qubit is not decoded for extended periods (which will occur if there are fewer decoders than qubits), then it can cause the decoder latency to increase due to an increase in undecoded errors.
@@
 Given the challenges in sharing decoder hardware and to understand how it can be enabled, we characterized representative FTQC workloads to understand the decoding requirements from a performance and reliability perspective.
-Our characterization using a lattice surgery compiler (watkins2023high, ) revealed that there is a limited amount of operational parallelism due to long sequences of $T$ and $H$ gates resulting from Clifford + $T$ decomposition, which is necessary for universality.
-This is crucial, as non-Clifford gates are the reason real-time decoding is necessary for FTQC.
+Our characterization using a lattice surgery compiler (Watkins et al., 2024) revealed that there is a limited amount of operational parallelism due to long sequences of $T$ and $H$ gates resulting from Clifford + $T$ decomposition, which is necessary for universality.
+This is crucial, as non-Clifford gates are the reason why real-time decoding is necessary for FTQC (Terhal, 2015).
 Non-Clifford operations are the only operations where the decoding is in the critical path, and fortunately, they occur in a highly serialized manner.
@@
 Furthermore, we design efficient decoder scheduling policies for such systems.
-Such a system can be visualized in Fig.
-1(b).
 We propose a scheduling policy that minimizes the Longest Undecoded Sequence, termed as the MLS policy.
-We compare it with the Round Robin (RR) and Most Frequently Decoded (MFD) policies – our evaluations show that the MLS policy can reduce the number of hardware decoders by up to 10$\times$ while ensuring that no logical qubit remains undecoded for a significantly long period.
-We also propose a noise-adaptive scheduling policy that can prioritize decoding of logical qubits that incur a sharp increase in the physical error rate due to phenomena such as cosmic rays (McEwen2021, ) and leakage due to heating (Miao2023, ).
-This involves a simple detector that can schedule decoding for a logical qubit in case the syndromes for that logical qubit show a sudden increase in bit-flips.
-Next, we show how some decoding tasks can be offloaded to software to further improve the efficacy of decoder scheduling policies.
-Balancing compute and memory is a classic architectural problem, and we use VQD to explore these trade-offs.
-Prior research on decoders and classical processing required for fault-tolerant quantum computers have focused on reducing the hardware cost of implementing decoders by making approximations in the decoding algorithm, sometimes at the cost of accuracy (Vittal2023, ; Riverlane2023, ; Alavisamani2024, ).
-With VaDER, we show that even if individual decoders have a high hardware cost, the overall cost can be reduced significantly by virtualizing decoders.
+We compare it with the Round Robin (RR) and Most Frequently Decoded (MFD) policies – our evaluations show that the MLS policy can reduce the number of hardware decoders by up to 10$\times$ for algorithmic logical qubits and 4$\times$ when magic state distillation factories are also considered, while ensuring that no logical qubit remains undecoded for a significantly long period.
+We show that offloading decoding tasks to software also reduce the longest period for which a logical qubit remains undecoded by up to 40%.
+We also propose a dynamic scheduling mechanism that can prioritize decoding for logical qubits affected by error bursts due to phenomena such as cosmic rays (McEwen et al., 2021) and leakage due to heating (Miao et al., 2023).
+We also analyze the effect of decoding latencies on the efficacy of virtualization.
+Our studies show that slow decoders are not amenable to virtualization, and QEC codes such as qLDPC codes will require significantly faster decoders than that are available today for virtualization to be possible.
+Through VQD, we show that the design space for decoders can be reduced from optimizing for latency, accuracy, and scalability (Figure 1(b)) to just latency and accuracy (Figure 1(c)).
+With VQD, we show that even if individual decoders have a high hardware cost, the overall cost can be reduced significantly by virtualizing decoders.
 2. Quantum Error Correction and Decoding
@@
 2.1. Quantum Error Correction
-Quantum Error Correction (QEC) improves the reliability of a system by utilizing many physical qubits to encode a single logical qubit (Shor1995, ; Knill1997, ).
-Most QEC codes can be categorized as stabilizer codes (Terhal2015, ) – some promising stabilizer codes include quantum Low Distance Parity Check (qLDPC) codes (Bravyi2024, ) and Surface Codes (Fowler2012, ).
+Quantum Error Correction (QEC) improves the reliability of a system by utilizing many physical qubits to encode a single logical qubit (Shor, 1995; Knill and Laflamme, 1997).
+Most QEC codes can be categorized as stabilizer codes (Terhal, 2015) – some promising stabilizer codes include quantum Low Distance Parity Check (qLDPC) codes (Bravyi et al., 2024) and Surface Codes (Fowler et al., 2012).
 Owing to their relatively relaxed connectivity requirements that can be realized with hardware available today, we focus specifically on the Surface Code.
-Note that this work can be extended to other QEC codes apart from the Surface Code.
 Figure 2. (a) A logical qubit ($d=3$); (b) Syndrome generation and measurements; (c) A typical procedure for detecting errors by decoding syndromes.
@@
 [some figure]
-A single rotated Surface Code patch of distance $d=3$ is shown in Fig.
-2(a) (Horsman2012, ).
+A single rotated Surface Code patch of distance $d=3$ is shown in Figure 2(a) (Horsman et al., 2012).
 Black circles denote data qubits and each data qubit is connected to $X$ and $Z$ parity qubits.
-The Surface Code works by repeatedly measuring syndromes, which correspond to the measurements performed on all $X$ and $Z$ measure qubits in a patch after performing the sequence of gates shown in Fig.
-2(b).
+The Surface Code works by repeatedly measuring syndromes, which correspond to the measurements performed on all $X$ and $Z$ measure qubits in a patch after performing the sequence of gates shown in Figure 2(b).
 By repeatedly measuring these syndromes, both bit-flip and phase-flip errors occurring within a patch can be detected by the decoder.
 2.2. Decoding Errors
-Fig.
-2(c) shows the general procedure of how any generic QEC code works – syndromes are constantly being generated, which are then fed to a decoder.
+Figure 2(c) shows the general procedure of how any generic QEC code works – syndromes are constantly being generated, which are then fed to a decoder.
 Syndromes contain information about which qubits have flipped in every round of syndrome measurements, and these flips allow the decoder to determine what errors on the data qubits caused those flips.
 Since errors can always be expressed in the form of Pauli gate, they can be corrected in software without executing any physical operations on the logical qubits.
-This is achieved by updating the Pauli frames for all data qubits that make up that logical qubit, which adjusts the interpretation of future measurements by accounting for the error that was detected (Terhal2015, ).
-For the Surface Code, the decoding problem is commonly formulated as a Minimum Weight Perfect Matching (MWPM) problem, which leverages a graph representation of the syndrome measurements (Fowler2015, ; Wu2022, ).
+This is achieved by updating the Pauli frames for all data qubits that make up that logical qubit, which adjusts the interpretation of future measurements by accounting for the error that was detected (Terhal, 2015).
+For the Surface Code, the decoding problem is commonly formulated as a Minimum Weight Perfect Matching (MWPM) problem (Fowler, 2015; Wu et al., 2022).
+2.3. Non-Clifford Gates
+On a Surface Code error-corrected quantum computer, all Clifford gates can be performed reliably either in software or via logical operations performed via Braiding (Fowler et al., 2012) or Lattice Surgery (Horsman et al., 2012).
+However, non-Clifford gates such as the $T$ gate cannot be applied in a fault-tolerant manner directly.
+This is because logical qubits initialized with the $T$ gate will have an error probability equal to the underlying physical error rate of the system, $p$, thus making them impure (Litinski, 2019b).
+However, multiple impure states can be used to distill fewer, purer logical qubits with a $T$ state (known as a magic state $|m\rangle$) – this process is known as magic state distillation (Bravyi and Haah, 2012; Gupta et al., 2024).
 Figure 3. Consumption of a magic ($T$) state with LS.
@@
 [some figure]
-2.3. Non-Clifford Gates
-On a Surface Code error-corrected quantum computer, all Clifford gates can be performed reliably either in software or via logical operations performed via Braiding (Fowler2012, ) or Lattice Surgery (Horsman2012, ).
-However, non-Clifford gates such as the $T$ gate cannot be applied in a fault-tolerant manner directly.
-This is because logical qubits initialized with the $T$ gate will have an error probability equal to the underlying physical error rate of the system, $p$, thus making them impure (Litinski2019magic, ).
-However, multiple impure states can be used to distill fewer, purer logical qubits with a $T$ state (known as a magic state $|m\rangle$) – this process is known as magic state distillation (Bravyi2012, ; Gupta2024, ).
-Fig.
-3 shows how a $T$ gate can be applied to a logical qubit $P$ in a fault-tolerant manner by using Lattice Surgery (LS) (Litinski2019, ).
+Figure 3 shows how a $T$ gate can be applied to a logical qubit $P$ in a fault-tolerant manner by using Lattice Surgery (LS) (Litinski, 2019a).
 The magic state $|m\rangle$ is a purified $T$-state.
-Since a non-Clifford gate is being applied, all prior errors that affected $P$ must be known before $|m\rangle$ is applied to prevent errors from spreading (Terhal2015, ).
-Lattice Surgery can be used to perform a $Z\otimes Z$ operation on $P$ and $|m\rangle$ to apply the magic state (Litinski2019, ; Litinski2019magic, ).
-Once the $Z\otimes Z$ operation is performed, the decoding result of $P$ prior to Lattice Surgery is combined with the decoding result of Lattice Surgery multi-body measurement and the measurement of the patch containing the magic state $|m\rangle$ to determine an appropriate Clifford correction [^1] The auto-corrected $\pi/8$ gate in (Litinski2019, ; Litinski2019magic, ) uses an additional ancillary qubit that has not been shown..
+Since a non-Clifford gate is being applied, all prior errors that affected $P$ must be known before $|m\rangle$ is applied to prevent errors from spreading (Terhal, 2015).
+Lattice Surgery can be used to perform a $Z\otimes Z$ operation on $P$ and $|m\rangle$ to apply the magic state (Litinski, 2019a, b).
+Once the $Z\otimes Z$ operation is performed, the decoding result of $P$ prior to Lattice Surgery is combined with the decoding result of Lattice Surgery multi-body measurement and the measurement of the patch containing the magic state $|m\rangle$ to determine an appropriate Clifford correction [^1] The auto-corrected $\pi/8$ gate in (Litinski, 2019a, b) uses an additional ancillary qubit that has not been shown..
 This correction needs to be known before the next logical operation involving a non-Clifford gate.
-[…]
-[some figure]
+We assume the use of auto-corrected Pauli-Product Rotations (PPR) (Litinski, 2019b) for all non-Clifford gates.
 Figure 4.
-(a) Exponential increase in the memory required to store undecoded syndromes for different number of available decoders for the wstate-60 benchmark – 200 decoders correspond to assigning a decoder to every logical qubit; (b) Number of $T$-gates required for different workloads; (c) Exponential increase in the decoder latency per round (in nanoseconds) as the number of rounds is increased from $d$ to $20d$ – the increase is higher for larger code distances ($p=10^{-4}$); (d) Slow increase in the logical error rate as the number of rounds of error correction increase with $p=10^{-3}$, $p=10^{-4}$; (e) Histogram of the number of concurrent critical decodes for different workloads with the EDPC layout – there is limited parallelism as far as critical decodes are concerned (f) Number of FPGAs required for decoders to implement different workloads with (i) a Single $T$-factory and (ii) Optimal number of $T$-factories.
+(a) Exponential increase in the decoder latency per round (in nanoseconds) as the number of rounds is increased from $d$ to $20d$ – the increase is higher for larger code distances ($p=10^{-4}$); (b) Slow increase in the logical error rate as the number of rounds of error correction increase with $p=10^{-3}$, $p=10^{-4}$; (c) Histogram of the number of concurrent critical decodes for different workloads with the EDPC layout – there is limited parallelism as far as critical decodes are concerned; (d) Estimated total decoders required for different workloads with the EDPC layout and 24 distillation factories for every workload.
 […]
@@
 2.4. Critical Decodes
-For a logical qubit that is only executing Clifford gates, errors can be decoded at any point of time (even after the experiment has ended).
-This is because all Clifford corrections can be commuted to the end of the circuit, essentially allowing the syndromes to be post-processed rather than decoded in real-time (Terhal2015, ).
+For a logical qubit that is only executing Clifford gates, errors can be decoded at any point in time (even after the experiment has ended).
+This is because all Clifford corrections can be commuted to the end of the circuit, essentially allowing the syndromes to be post-processed rather than decoded in real-time (Terhal, 2015).
 However, universal fault-tolerant quantum computer require non-Clifford gates such as the $T$-gate – this makes real-time decoding a necessity since syndromes for a patch must be decoded before the application of a non-Clifford gate.
-We call decodes that must happen before the application of a non-Clifford gate critical decodes since all syndromes generated up to that point for that logical qubit must be decoded before computation can proceed.
+We call decodes that must happen before the application of a non-Clifford gate critical decodes, since all syndromes generated up to that point for that logical qubit must be decoded before computation can proceed.
+Finally, performing non-Clifford operations on logical qubits encoded in quantum Low Distance Parity Check (qLDPC) codes require a critical decode to measure the ancillary system required to transfer the logical qubit to a Surface Code patch (Cross et al., 2024).
 3. Classical Processing Requirements
-Having explained the relevant details about QEC and the role of decoders and critical decodes, we now cover the classical processing requirements for FTQC.
+We now discuss the classical processing requirements for FTQC.
 3.1. Syndrome Generation and Processing Rates
 As discussed in Section 2, applying non-Clifford gates requires decoders to be up-to date with the latest syndrome for a logical qubit before computation can proceed.
-As shown by Terhal (Terhal2015, , p.
-20), if the rate at which syndromes are generated $r_{gen}$ is faster than the rate at which they are processed $r_{proc}$, the memory required for storing syndromes that are yet to be decoded increases exponentially (referred to as the backlog problem).
-This exponential increase in the memory required also leads to an exponential increase in the runtime of the workload.
-Fig.
-4(a) shows the exponential increase in the memory required to store undecoded syndromes for the wstate-60 benchmark with the number of rounds for various numbers of available decoders.
-For larger and longer running workloads, the memory requirements will be much higher.
-By reducing the number of available decoders, $r_{proc}$ is effectively reduced, leading to the exponential growth in memory.
-Decoding is necessary at runtime when consuming $T$-states, as discussed in Section 2, and Fig.
-4(b) shows that it will be a frequent operation for most workloads.
-$r_{proc}$ must be consistently greater than $r_{gen}$ to prevent an exponential increase in memory requirements.
-3.2. Why is Real-Time Decoding Needed?
-[…]
-Leaving syndromes undecoded for even tens of rounds can result in an exponential slowdown in the decoder processing rate.
+As shown by Terhal (Terhal, 2015, p.
+20), if the rate at which syndromes are generated $r_{gen}$ is faster than the rate at which they are processed $r_{proc}$, the time required to execute a workload increases exponentially with the number of non-Clifford $T$ states required by the program (the backlog problem (Terhal, 2015)).
+Now, if every decoder in the system has an $r_{proc}>r_{gen}$, and every logical qubit has its own decoder, there will be no backlog of syndromes to be decoded and hence no slowdown in the system.
+Since non-Clifford operations will be very frequent in FTQC systems (Beverland et al., 2022b), ensuring that $r_{proc}$ is consistently greater than $r_{gen}$ is a critical requirement for FTQC systems.
+3.2. Why is Real-Time Decoding Necessary?
+This requirement for syndromes to be processed faster than they are generated has motivated research to build fast and accurate hardware decoders for the Surface Code (Smith et al., 2023; Ravi et al., 2023; Barber et al., 2023; Vittal et al., 2023; Alavisamani et al., 2024), especially for systems using superconducting qubit architectures due to their fast gate times.
+While the syndrome processing rates achieved by these decoders are far higher than typical syndrome generation rates achieved today (Acharya et al., 2023; Bluvstein et al., 2023), leaving syndromes undecoded for many successive rounds can be problematic.
+Figure 4(a) shows how the decoder latency normalized to the number of rounds (latency per round) processed by the decoder can increase exponentially with the number of rounds of undecoded syndromes, especially for $d=7,9$ (circuit-level noise $p=10^{-4}$).
+The number of rounds was chosen as a multiple of $d$ since it represents the shortest period required for executing a logical operation (Horsman et al., 2012).
+This slowdown is easily explainable by the fact that errors will accumulate the longer a logical qubit remains undecoded, thus requiring more corrections to be performed.
+This can result in a significant slowdown, and hence a decrease in $r_{proc}$ leading to higher memory requirements to store undecoded syndromes.
+However, note that the slowdown in $r_{proc}$ will result in more rounds of error correction required to complete the computation.
+Figure 4(b) shows how the logical error rate grows slowly with the number of rounds for $p=10^{-3}$, $p=10^{-4}$ respectively.
+Code distances are selected to achieve a target logical error rate after the $N$ rounds it takes to complete a program (Beverland et al., 2022b; Blunt et al., 2024) – critical decodes delayed exponentially due to a slow $r_{proc}$ will exacerbate the logical error rate, since more rounds would be needed to complete the computation.
 3.3. Concurrency and Delayed Decoding
@@
 If there are many concurrent critical decodes that occur frequently during the execution of a program, an appropriate number of decoders will be needed to process these critical decodes.
-The level of concurrency depends entirely on the layout used to build the quantum computer – layouts such as the Compact (Litinski2019, , p.
-7), Fast (Litinski2019, , p.
-9), and the Edge-Disjoint Path (EDPC) (Beverland2022edpc, ) are some proposed layouts that can be used to build fault-tolerant quantum computers using the Surface Code.
+The level of concurrency depends entirely on the layout used to build the quantum computer – layouts such as the Compact (Litinski, 2019a, p.
+7), Fast (Litinski, 2019a, p.
+9), and the Edge-Disjoint Path (EDPC) compilation (Beverland et al., 2022a) are some proposed layouts that can be used to build fault-tolerant quantum computers using the Surface Code.
 Layouts determine the number of physical qubits required – for example, the Compact layout requires the fewest physical qubits since there is a single routing lane at the cost of completely serializing operations.
@@
 What is the average level of concurrency when executing a quantum program on a fault-tolerant quantum computer?
-Fig.
-4(e) shows a histogram of the number of critical decodes [^3] For the Surface Code, we assume every logical qubit requires two decoders – one each for the $X$ and $Z$ observables.
-for select workloads generated by the Lattice Surgery Compiler (watkins2023high, ).
+Figure 4(c) shows a histogram of the number of critical decodes [^2] For the Surface Code, we assume every logical qubit requires two decoders – one each for the $X$ and $Z$ observables.
+for select workloads generated by the Lattice Surgery Compiler (Watkins et al., 2024).
 This histogram shows that the peak concurrency is attained very infrequently.
 This implies that most logical qubits function as memory qubits or execute Clifford gates more often than $T$-gates, and this can allow some qubits to not be decoded in real-time.
-Quantum programs are serial in terms of $T$-gates applied – not every qubit always requires access to a fast decoder.
+Quantum programs are serial in terms of $T$-gates applied – not every qubit requires constant access to a fast decoder.
 3.4. Goal: Make Classical Processing Efficient
-Provisioning a hardware decoder for every logical qubit in the system can be resource intensive – Fig.
-4(f) shows an estimate of the number of FPGAs required just for decoding when (i) a single distillation factory is used, and (ii) an optimal number of distillation factories are used for different workloads ($\sim$10FPGA LUTs/decoder (Vittal2023, ; Riverlane2023, )).
-This estimation was done using the Azure QRE (Beverland2022, ) that uses the Fast layout.
+Provisioning a hardware decoder for every logical qubit in the system can be resource intensive.
+Figure 4(d) shows an estimate of the total number of decoders required for all logical qubits (=algorithmic, ancillary, magic state storage, and magic state distillation logical qubits) when using the EDPC layout for different workloads.
+These estimates were obtained by using Lattice Surgery Compiler (Leblond et al., 2024; Watkins et al., 2024).
 Note that the total hardware requirement will be significantly higher because of control and readout components.
 Having shown how the syndrome processing rate is crucial in ensuring that computation does not require excessive memory and time and how quantum programs are inherently serial in terms of critical decodes, the question we seek to answer in this work is–
-How can we minimize the use of hardware decoders and lower classical processing costs without sacrificing the performance and reliability?
-4. Virtual Quantum Decoders
-We now show how the number of hardware decoders can be reduced to be less than the number of logical qubits in the system, and how decoding can be scheduled to prevent excessive accumulation of undecoded syndromes.
-Figure 5. (Left) Decoders for every logical qubit; (Right) Time-division multiplexing of decoders between qubits.
-[…]
-[some figure]
+How can we minimize the use of hardware decoders and lower classical processing costs without sacrificing performance and reliability?
+4. VQD: Virtual Quantum Decoding
+We now discuss decoder virtualization, where there are fewer hardware decoders than logical qubits in the system, and how decoding can be scheduled optimally.
 4.1. Working with Fewer Hardware Decoders
 Reducing the number of available decoders implies that qubits will share hardware resources, resulting in time-division multiplexing of decoder instances among logical qubits.
-Fig.
-5 shows how compared to a system with decoders for every logical qubit in the system ($N$ qubits, $N$ decoders), a system with fewer ($M$, $N>M$) decoders will require resources to be shared with time.
+Figure 5 shows how compared to a system with decoders for every logical qubit in the system ($N$ qubits, $N$ decoders), a system with fewer ($M$, $N>M$) decoders will require resources to be shared with time.
 Time-division multiplexing of hardware resources will require the following considerations: (i) If the number of critical decodes at a given time step exceed the number of hardware decoders, the overflowing critical decodes will have to be deferred to the next available time step, and (ii) Qubits cannot be left undecoded for extended periods of time.
@@
 Before we discuss different decoder scheduling policies, it is important to understand the time granularity at which any scheduling policy will operate on.
-Since logical operations (Clifford or non-Clifford) in a Surface Code error-corrected quantum computer will require at least $d$ rounds before the next operations, we define a slice (watkins2023high, ) as the smallest time step between logical operations that a decoder scheduling policy can work on.
+Since logical operations (Clifford or non-Clifford) in a Surface Code error-corrected quantum computer will require at least $d$ rounds before the next operations, we define a slice (Watkins et al., 2024) as the smallest time step between logical operations that a decoder scheduling policy can work on.
 Every slice consists of $d$ rounds of syndrome measurements, thus making the scheduling policy agnostic of the actual code distance used.
-Figure 6.
-(a) Illustration of the longest undecoded sequence – $Q_{1}$ has the longest undecoded sequence before the last decode; (b) MFD policy – undecoded qubits are sorted according to the number of critical decodes they are involved in after time slice $t$ and $M-C$ qubits are selected from this sorted list; (c) RR policy – the $M-C$ qubits decoded in time slice $t$ are not decoded in time slice $t+1$; (d) MLS policy – undecoded qubits are sorted according to their undecoded sequence lengths and $M-C$ qubits are selected from this sorted list.
+Figure 5. (Left) Decoders for every logical qubit; (Right) Time-division multiplexing of decoders between qubits.
 […]
@@
 Since most quantum programs do not have any control-flow instructions, scheduling can be performed statically.
-Scheduling decoders is similar to CPU scheduling performed by all operating systems today, where the number of processes is more than the number of available processor cores (Liu1973, ).
+Scheduling decoders is similar to CPU scheduling performed by all operating systems today, where the number of processes is more than the number of available processor cores (Liu and Layland, 1973).
 Longest Undecoded Sequence: To quantify the fairness of a decoder scheduling policy, we use ‘Longest Undecoded Sequence’, which measures how well the decoders are servicing all logical qubits.
 A large undecoded sequence length implies that a qubit has been left undecoded for a long time – increasing the memory consumed to store undecoded syndromes.
-Fig.
-6(a) shows an example of determining the longest undecoded sequence length.
+Figure 6(a) shows an example of determining the longest undecoded sequence length.
 Consider an arbitrary time slice $t$ in the execution of a quantum program.
@@
 The second will assign all the remaining $M-C$ hardware decoders to the $N-C$ qubits based on the scheduling policy used.
-We now discuss three decoder scheduling policies (all policies are illustrated in Fig.
-6(b) – Fig.
-6(d)):
+We now discuss three decoder scheduling policies (all policies are illustrated in Figure 6(b) – Figure 6(d)):
+Figure 6.
+(a) Illustration of the longest undecoded sequence – $Q_{1}$ has the longest undecoded sequence before the last decode; (b) MFD policy – undecoded qubits are sorted according to the number of critical decodes they are involved in after time slice $t$ and $M-C$ qubits are selected from this sorted list; (c) RR policy – the $M-C$ qubits decoded in time slice $t$ are not decoded in time slice $t+1$; (d) MLS policy – undecoded qubits are sorted according to their undecoded sequence lengths and $M-C$ qubits are selected from this sorted list.
+[…]
+[some figure]
 4.2.1. Most Frequently Critically Decoded (MFD)
 A logical qubit that consumes a significant number of $T$-states during the execution of a program would have a frequent requirement of critical decodes – leaving such a logical qubit undecoded for more than a few slices would make subsequent critical decodes take longer, thus slowing down computation.
-This motivates the MFD scheduling policy that prioritizes decoding of logical qubits that have numerous critical decodes in the future at any given time slice.
+This motivates the MFD scheduling policy that prioritizes decoding of logical qubits that have numerous critical decodes in the future at any given time slice (Figure 6(b)).
 The MFD policy will ensure that future critical decodes have a minimized number of undecoded syndromes for the qubits that have frequent critical decodes.
@@
 4.2.2. Round Robin (RR)
-Derived from CPU scheduling policies used by operating systems, the RR policy does not prioritize any specific logical qubits – rather, it chooses $M-C$ qubits in a round-robin manner in every time slice to ensure fairness for all qubits in the system.
+Derived from CPU scheduling policies used by operating systems, the RR policy does not prioritize any specific logical qubits – rather, it chooses $M-C$ qubits in a round-robin manner (Figure 6(c)) in every time slice to ensure fairness for all qubits in the system.
 Caveats: For regions of a program where there are many critical decodes, the RR policy could still starve some logical qubits since $M-C$ will be much smaller, yielding a smaller window for decoders to be assigned.
@@
 The longest undecoded sequence length at any given time slice is an indicator of how well the decoder scheduling policy is servicing all qubits in the system.
-We use this as a motivator for the MLS policy, which tries to minimize the longest undecoded sequence at every time slice.
+We use this as a motivator for the MLS policy, which tries to minimize the longest undecoded sequence at every time slice (Figure 6(d)).
 The MLS policy works as follows: at any time slice $t$, qubits are sorted on the basis of their current undecoded sequence lengths.
@@
 4.3. Noise-Adaptive Decoder Scheduling
-While control-flow instructions would necessitate runtime scheduling of decoders, events such as cosmic rays (McEwen2021, ) and leakage (Miao2023, ) can result in a temporary burst of errors for some physical qubits in the lattice that can impact some logical qubits.
-Scheduling after a control-flow instruction can be performed using any static scheduling policies for the program after the control-flow instruction.
-However, since the static scheduling policies do not account for the error-rate, spikes in errors due to cosmic rays and leakage cannot be factored at runtime without hardware support.
 Figure 7.
@@
 [fig]
+While control-flow instructions would necessitate runtime scheduling of decoders, events such as cosmic rays (McEwen et al., 2021) and leakage (Miao et al., 2023) can result in a temporary burst of errors for some physical qubits in the lattice that can impact some logical qubits.
 Detecting a spike in the physical error rate will either require errors to be decoded or additional hardware modules to detect the spike.
-While error-correcting codes can tolerate temporary increases in the error-rate (Google2023, ), the increase in errors can result in longer decoding latencies since the decoding task becomes harder with more errors.
-If a logical qubit affected by these events is not scheduled for decoding immediately after the event, decoding it before applying a non-Clifford gate could take longer, thus delaying the operation and causing a slowdown.
-As shown in Fig.
-7(a), an increase in the physical error rate results in a higher number of bit-flips (especially for larger code distances), which can be detected with simple components in the control hardware.
-Fig.
-7(b) shows how additional flips can be detected and used to dynamically prioritize the decoding of an arbitrary patch $P$, which suffers from a temporary burst of errors.
+As shown in Figure 7(a), an increase in the physical error rate results in a higher number of bit-flips (especially for larger code distances), which can be detected with simple components in the control hardware.
+Figure 7(b) shows how additional flips can be detected and used to dynamically prioritize the decoding of an arbitrary patch $P$, which suffers from a temporary burst of errors.
 Note that the detection is different from decoding – we are merely predicting that there are more errors due to higher bit-flips in the syndromes.
+Unlike prior work (Suzuki et al., 2022), this mechanism only prioritizes the decoding of logical qubits affected by error bursts in a virtualized decoder environment.
+However, this prioritization can result in reduced efficacy of virtualization, due to fewer decoders being available for scheduling during error burst events.
+We evaluate the overheads of such error bursts for virtualization in Section 7.
 4.4. Offloading to Software Decoders
-Software decoders are slow and also have a higher variance in decoding latencies (Delfosse2023, ).
+Software decoders are slow and also have a higher variance in decoding latencies (Delfosse et al., 2023).
 However, when scheduling decoding tasks for logical qubits, software decoders can be leveraged to further reduce the undecoded sequence lengths.
-As shown in Fig.
-7(c), some syndromes for a logical qubit can be offloaded to software while the hardware decoders are busy elsewhere.
+As shown in Figure 7(c), some syndromes for a logical qubit can be offloaded to software while the hardware decoders are busy elsewhere.
 To prevent scheduled hardware decoding from being delayed, a buffer (three slices in this example) must be used to ensure that the software offloading completes before the next hardware decode.
-5. Decoding for Distillation Factories
+4.5. Decoding for Distillation Factories
 The decoder scheduling policies in the previous sections catered only to the decoding of algorithmic logical qubits (data logical qubits, magic state storage, ancillary logical qubits required for Lattice Surgery).
-In this section, we discuss decoding for distillation factories.
-5.1. Distillation Factories
 Magic state distillation factories generate few low-error logical qubits with non-Clifford states from many high-error logical qubits.
-Distillation factories run for very short periods at a time – this allows for smaller code distances to be used for creating the logical qubits for distillation (Litinski2019magic, ).
-As shown in Fig.
-8, the error probability of a magic state is low enough to be useful even with $d=7$ ($d$ refers to $d_{X}$ in (Litinski2019magic, )).
-Figure 8. Output error probabilities of magic states after 15-1 distillation for different $d$ (Litinski2019magic, ) ($p=10^{-4}$).
+Distillation factories run for very short periods at a time – this allows for smaller code distances to be used for creating the logical qubits for distillation (Litinski, 2019b).
+Smaller code distances for distillation factories provide two main benefits: the number of physical qubits required are much lower, and more importantly, both hardware (Vittal et al., 2023; Smith et al., 2023; Delfosse, 2020; Caune et al., 2023) and software (Delfosse et al., 2023) decoders are faster and less complex.
+For example, LUT based decoders (Das et al., 2022a) have been shown to be effective up to $d=5$ without requiring significant hardware resources.
+Predecoders (Delfosse, 2020; Ravi et al., 2023; Smith et al., 2023) reduce the complexity and decoding effort required for lower code distances as well.
+Compared to algorithmic logical qubits, which require a large code distance to survive millions of error correction cycles (Beverland et al., 2022b; Blunt et al., 2024), the decoding requirements of distillation factories are far more relaxed, which reduces the hardware resource requirements as well.
+Magic state distillation has lower hardware requirements for decoding due to lower code distances involved – and virtualization can further reduce them.
+4.6. Impact of Decoding Latencies
+So far, we have considered decoder virtualization without considering the decoder latency.
+However, the former is heavily impacted by the latter.
+Consider a scenario where a logical qubit has $R$ undecoded rounds prior to a critical/scheduled decode.
+Now, the decoder will take a finite time to process these $R$ rounds worth of syndromes based on its per round decoding latency $T_{proc}$.
+However, during the time the initial $R$ rounds of syndromes are processed, syndromes are continuously generated based on the syndrome generation time $T_{gen}$.
+Thus, at any time $t$, the total rounds processed are $P(t)=t/T_{proc}$, the syndrome rounds generated are $G(t)=t/T_{gen}$, and so, the outstanding rounds to be processed can be defined as $S(t)=R+G(t)-P(t)$.
+The total time taken for the decoder to catch up is thus $t_{total}$, with $S(t_{total})=0$, and hence:
+(1) $t_{total}=\frac{R}{T_{proc}^{-1}-T_{gen}^{-1}};(iff~{}~{}T_{proc}<T_{gen})$
+Figure 8.
+(a) Slowdown in the processing of outstanding syndromes for different normalized decoder latencies $t_{D}=T_{proc}/T_{gen}$.
+Slower decoders require more time (and thus more rounds) to catch up with the present syndrome; (b) Differences in the total decoding task for different initial undecoded rounds – a $t_{D}=0.99$ increases the number of rounds to be decoded by a 100$\times$.
 […]
-[fig]
-5.2. Using Fast, Low-Footprint Decoders
-Smaller code distances for distillation factories provide two main benefits: the number of physical qubits required are much lower, and more importantly, both hardware (Vittal2023, ; Smith2023, ; Delfosse2020, ; Caune2023, ) and software (Delfosse2023, ) decoders are faster and less complex.
-For example, LUT based decoders (Das2022lilliput, ) have been shown to be effective up to $d=5$ without requiring significant hardware resources.
-Predecoders (Delfosse2020, ; Ravi2023, ; Smith2023, ) reduce the complexity and decoding effort required for lower code distances as well.
-Compared to algorithmic logical qubits, which require a large code distance to survive millions of error correction cycles (Beverland2022, ; Blunt2024, ), the decoding requirements of distillation factories are far more relaxed, which reduces the hardware resource requirements as well.
-The decoding overhead of magic state distillation is significantly lower than algorithmic logical qubits – it can thus leverage lightweight decoders, considerably reducing the hardware cost for distillation factories.
+[some figure]
+$T_{proc}$ has a considerable effect on the efficacy of decoder virtualization, since a slower processing rate would mean that a huge number of additional rounds will be required for the decoder to be able to catch up to the present syndrome being generated.
+This can be seen from Figure 8(a), where the decoder latency is normalized by the syndrome generation time $t_{D}=T_{proc}/T_{gen}$ and the slowdown is defined as the ratio of the total time taken to process $R$ undecoded rounds (Equation 1) and the ideal time when extra syndromes are added to the decoding task ($t_{total}^{\prime}=R/T_{proc},slowdown=t_{total}/t_{total}^{\prime}$).
+This slowdown is agnostic of the actual implementation of the decoder – techniques like parallel window decoding (Skoric et al., 2023) will only help reduce the value of $t_{D}$.
+Figure 8(a) shows the total decoding task ($P(t_{total})$), which represents the total syndromes that need to be processed for an initial $R$ undecoded rounds.
+From Figure 8(a), we see that the closer the decoding processing rate is to the syndrome generation rate ($t_{D}=0.99$), the harder virtualization will be, since even $R=10$ will require $10^{4}$ rounds before the decoder can catch up.
+This is a crucial limitation for decoder virtualization, and it also suggests that faster, more accurate decoders will be more desirable than fast decoders which have a smaller hardware footprint at the cost of accuracy.
+Figure 8(b) shows an example for cases where there are 10/100 undecoded rounds before the decoder is invoked for a logical qubit – a slower decoder will require thousands of extra rounds to catch up compared to a faster decoder!
+Running so many additional rounds will slow down the system, reduce the error budget per round of computation, exacerbating the logical error rates.
+In Section 7, we will show how slow decoders associated with qLPDC codes are not amenable for virtualization, due to the slowdown they incur for even modest undecoded sequence lengths.
+This represents a limit of virtualization, and motivates further research in developing fast, accurate decoders for all QEC codes.
+Figure 9. System architecture for virtualized decoding.
+[…]
+[some figure]
+5. System Architecture with Virtual Decoders
+Enabling the virtualization of decoders in an FTQC system requires (i) memory to store syndromes of logical qubits that are not being decoded at any time, and (ii) a mechanism for selecting the qubits to be decoded by a specific decoder at any time.
+Figure 9 shows the system architecture that can be used for enabled virtualized decoders.
+The Arbiter is a key component that routes syndromes from the readout system to the decoders.
+Since there will be fewer decoders than logical qubits, syndromes will be routed based on the static decoding schedule prepared during compile time.
+To allow decoders to access the syndromes of a specific logical qubit at any given point in time, the arbiter attaches a tag to the syndromes before routing them to the appropriate decoder.
+This tag quickly identifies the logical qubit to which a syndrome belongs to.
+Every decoder consists of a syndrome memory, and the decoding schedule can be used by the control software to access the syndromes of any logical qubit from the syndrome memory.
+Software offloading can be easily integrated in the design by enabling the syndrome memory to be read from the control software.
+Note that this will require high-speed links between the control software and the decoder modules, such as the low-latency ethernet used in recent experimental demonstrations (Acharya et al., 2024).
+The control software can specify the tag and the syndromes to be read from the syndrome memory.
+This will however require fusing the software decoding results with the hardware decoding results.
+Fusion of separate decoding tasks has also been demonstrated recently by Google (Acharya et al., 2024).
+To enable dynamic decoder scheduling, the control software need only trigger decoding for the necessary logical qubit(s) by amending the decoding schedule.
+Since this could trigger the decoding of any arbitrary logical qubit, the syndrome memories must support random accesses.
+No other changes to the Arbiter are necessary, since syndromes can be routed to the same memory regardless of dynamic scheduling that could occur later in time.
+Figure 10. Compilation framework – the rewrite pass changes the IR schedule based on the number of available decoders.
+[…]
+[some figure]
 6. Methodology
-We now describe the methodology used to evaluate different decoder scheduling policies and for estimating classical resources required for executing workloads on a Surface Code error-corrected quantum computer.
+We now describe the methodology used to evaluate different decoder scheduling policies and classical processing requirements.
 6.1. Compiler
-We use the Lattice Surgery Compiler (LSC) (watkins2023high, ) to generate Intermediate Representations (IR) of workloads that can be executed on an error-corrected quantum computer using the Surface Code with lattice Surgery.
-LSC can generate IR that denote Lattice Surgery instructions from the QASM (qasm, ) representation of a workload.
-LSC handles mapping and routing based on the layout provided to the compiler.
+We use the Lattice Surgery Compiler (LSC) (Watkins et al., 2024) to generate Lattice Surgery Instructions (LLI) Intermediate Representations (IR) of workloads that can be executed on an error-corrected quantum computer using the Surface Code with lattice Surgery.
+LSC can generate IR that denote Lattice Surgery instructions from the QASM (Cross et al., 2017) representation of a workload.
+The Lattice Surgery instructions generated by LSC are a combination of Clifford operations and multi-body measurements used for implementing Pauli-Product Rotations (PPR) (Litinski, 2019a).
+LSC handles mapping and routing based on the layout provided to the compiler, as shown in Figure 10.
 We configure LSC to use a ‘wave’ scheduling that maximizes the number of concurrent instructions executed in every time slice.
-LSC also uses Gridsynth (gridsynth, ) to deal with arbitrary rotations.
+LSC uses Gridsynth (Ross and Selinger, 2014) to deal with arbitrary rotations.
 However, since it is still under development, LSC has some limitations:
   • LSC is limited to multi-body measurements between only two logical qubits.
-  • LSC abstracts away distillation factories, only magic state storage sites are considered.
+  • LSC treats distillation factories as black boxes.
   • LSC works for a limited set of layouts and is extremely slow for large workloads like shor.
 6.2. Simulation Framework
-Using the IR generated by LSC, we build a framework that can parse the IR and determine the critical decodes in every slice, generate a timeline of all operations, and assign decoders to all logical qubits depending on the scheduling policy.
-In case the number of critical decodes in a particular slice are more than the number of hardware decoders configured, decoder-resources can rewrite the IR to defer critical decodes to the next slice (potentially increasing the execution time of the program).
-Layouts: We use three layouts for our evaluations – Fast and Compact layouts (Litinski2019, ), and the EDPC layout (Beverland2022edpc, ).
+Using the IR generated by LSC, we build a framework [^3] https://anonymous.4open.science/r/decoder-resources-5EC4/ that can parse the IR and determine the critical decodes in every slice, generate a timeline of all operations, and assign decoders to all logical qubits depending on the scheduling policy.
+In case the number of critical decodes in a particular slice are more than the number of hardware decoders configured, decoder-resources can rewrite the IR to defer critical decodes to the next slice.
+Layouts: We use three layouts for our evaluations – Fast, Compact (Litinski, 2019a), and the EDPC (Beverland et al., 2022a) layouts.
 The Compact layout uses the fewest logical qubits and, due to a single routing lane, allows only one magic state to be consumed per time slice – we thus use it only to compare total execution times with the Fast and EDPC layouts.
-Benchmarks: We use benchmarks from MQT Bench (quetschlich2023mqtbench, ) and QASMBench (Li2023, ).
-We use shor-15, a chemistry workload gndstate-14, a NISQ workload qaoa-14444Used for its arbitrary rotations and similarity to chemistry workloads., random, wstate, and arithmetic workloads adder-28, multiplier-45, Quantum Fourier Transform qft-20 – which can be used as building blocks for other algorithms.
-Other Software: Stim (Gidney2021, ) was used for simulating stabilizer circuits to generate syndromes and error rates. Azure QRE (Beverland2022, ) was used for resource estimations.
+Benchmarks: We use benchmarks from (Quetschlich et al., 2023) and (Li et al., 2023).
+We use shor-15, phase estimation qpe-50, Quantum Fourier Transform qft-20, arithmetic workloads adder-28, multiplier-45.
+We also use a neural network workload dnn-51, W state workloads wstate-40/60, random unitary random-40, and an Ising model ising-66.
+Other Software: Stim (Gidney, 2021) was used for simulating stabilizer circuits to generate syndromes and error rates. Azure QRE (Beverland et al., 2022b) was used for resource estimations.
 7. Evaluations
 In this section, we present some results for different scheduling policies and savings in decoder hardware.
-Figure 9.
-(a) Max.
-concurrent critical decodes for all layouts and workloads; (b) Total execution time (in code cycles) needed to execute all workloads with different layouts (c) Estimated number of logical qubits for all workloads (code distances estimated by Azure QRE (Beverland2022, ) also annotated).
+(a) Max. concurrent critical decodes.
+(b) Total execution time (in code cycles)
+(c) Estimated number of logical qubits (code distances estimated by Azure QRE (Beverland et al., 2022b) also annotated).
+Figure 11. Baseline statistics.
 […]
@@
   2. (2)
-    How long do qubits go undecoded when using different scheduling policies?
+    How long do qubits go undecoded when using virtualization?
   3. (3)
-    How do scheduling policies affect memory usage when using virtualized decoders?
+    How does virtualization affect Magic State Distillation?
+  4. (4)
+    How do error bursts affect decoder resource requirements?
+  5. (5)
+    How do slow decoders for qLDPC codes affect the efficacy of virtualization?
 7.2. Baseline Statistics
-[…]
-[…]
+We consider the baseline to have decoders for every logical qubit in the system.
+Figure 11(a) shows the maximum number of critical decodes that occur during the execution of different workloads for all selected layouts.
+The Compact layout has a maximum of one critical decode per slice between two logical qubits.
+Figure 11(b) shows the total time required to finish all workloads444shor-15 and qpe-50 were stopped prematurely due to long runtimes (shor-15 after 100,000 time slices for all layouts, qpe-50 after 72,000 time slices for Fast, EDPC, and 100,000 time slices for Compact).
+– the EDPC and Fast layouts are significantly faster than Compact, and the EDPC layout has a very slight advantage over Fast for most workloads and also uses fewer qubits 11(c) (only algorithmic logical qubits).
+Due to its higher degree of parallelism (due to its better use of routing lanes), we only consider the EDPC layout for all further evaluations.
 7.3. Decoder Scheduling Efficacy
-Figure 10. Hardware decoders used for different configurations (EDPC layout).
+Figure 12 shows the three decoder configurations used in this work.
+  • The All Qubits configuration denotes the baseline where all qubits have a decoder.
+  • Max. Concurrency is the configuration where the number of hardware decoders in the system corresponds to the peak concurrent critical decodes for every workload shown in Figure 11(a).
+  • Midpoint refers to a configuration where the number of hardware decoders is the midpoint between the max. and min. concurrent critical decodes ($=\frac{Max.+Min.}{2}$).
+Figure 12. Hardware decoders normalized with the Midpoint configuration (EDPC layout). Decoders used by the Midpoint configuration are annotated below the benchmarks.
 […]
@@
 [some figure]
-Fig. 10 shows the three decoder configurations selected for this work.
-  • The All Qubits configuration denotes the baseline where all qubits have a decoder.
-  • Max. Concurrency refers to the configuration where the number of hardware decoders in the system corresponds to the peak concurrent critical decodes for every workload shown in Fig. 9(a).
-  • Midpoint refers to a configuration where the number of hardware decoders is the midpoint between the max. and min. concurrent critical decodes ($=\frac{Max.+Min.}{2}$).
-The minimum concurrent critical decodes corresponds to two critical decodes between two logical qubits.
-All evaluations are for algorithmic logical qubits, logical qubits required for distillation are not considered.
+The minimum concurrent critical decodes corresponds to one critical decode between two logical qubits because a multi-body measurement merges the two patches into a single decoding task.
+We first focus our evaluations on algorithmic logical qubits, and then incorporate distillation factories as well.
 For the All Qubits configuration, the longest undecoded sequence length will be zero, since every logical qubit has an assigned decoder.
@@
 Since these configurations use far fewer hardware decoders than qubits, the longest undecoded sequence is a good measure of whether qubits are being starved of decoding.
-Fig.
-11(b) shows the longest undecoded sequence lengths for the (a) Max.
+Figure 13(b) shows the longest undecoded sequence lengths for the (a) Max.
 Concurrency and (b) Midpoint configurations.
 The MFD policy leads to qubits being starved of decoding, since it prioritizes qubits that have frequent critical decodes.
-For the Max.
-Concurrency configuration, the gndstate and qaoa workloads do relatively well with the MFD configuration signifying that almost all logical qubits have a similar number of critical decodes, leading to a fairer scheduling.
-While the RR policy performs significantly better than the MFD policy, the MLS policy consistently performs better than both policies – MLS reduces the longest undecoded sequence lengths for almost every workload to $\sim 10$ slices for both Max.
+While the RR policy performs significantly better than the MFD policy, the MLS policy consistently performs better than both policies – MLS reduces the longest undecoded sequence lengths for almost every workload to $\sim$10 time slices for both Max.
 Concurrency and Midpoint configurations.
-(a) Max. Concurrency
-(b) Midpoint
-Figure 12. Peak memory required for storing undecoded syndromes for different scheduling policies when using the (a) Max. Concurrency; (b) Midpoint configurations.
+Figure 14. Memory required (normalized by the MLS policy) for undecoded syndromes with the Midpoint configuration.
 […]
@@
 The reduction in the longest undecoded sequence lengths also corresponds to lower memory usage for storing undecoded syndromes.
-This is crucial since reducing the number of hardware decoders will require more memory to store syndromes for qubits that have not been decoded.
-Fig.
-12(b) shows the memory required for different workloads with the (a) Max.
-Concurrency and (b) Midpoint configurations (the Azure QRE estimated the code distances used to determine the memory requirements).
-Due to longer undecoded sequences, the MFD policy can require up to 100 GB of memory for some workloads while the MLS policy rarely requires more than 100 MB of memory, which is orders of magnitude better than the MFD policy and 2-4x better than the RR policy.
-Figure 13. Total slices required by workloads with different decoder configurations normalized to the slices required by the baseline (MLS policy).
-[…]
-[some figure]
-7.3.3. Slowdown due to Fewer Decoders
-Fewer hardware decoders imply that some critical decodes in an arbitrary slice have to be deferred to subsequent slices, resulting in potentially more slices for completing the program.
-Fig.
-13 shows the number of slices required for all workloads normalized with respect to the baseline number of slices shown in Fig.
-9(b) for the Min.
-Concurrency, Max.
-Concurrency, and Midpoint configurations.
-Since the Min.
-Concurrency configuration allocates only four decoders for two critical decodes per slice, some workloads are slowed down by ¿10%.
-The Midpoint configuration however does not cause any slowdown except in the gndstate workload, and there is no slowdown for the Max.
-Concurrency configuration.
-Figure 14. Increase in the LER (lower is better) with the Midpoint configuration.
-[…]
-[some figure]
-7.3.4. Impact on Logical Error Rate
-As discussed in Section 3, delaying decoding does not affect the logical error rate (LER) by itself.
-However, leaving a qubit undecoded for a long time before a critical decode can slow the decoder down, thus requiring additional rounds to complete the computation.
-In this evaluation, we would thus like to show how the longest undecoded sequence length can impact the LER.
-Fig.
-14 shows how the MFD policy can increase the final LER of the computation due to longer undecoded sequence lengths.
-The RR policy increases the LER slightly for adder-28 and random-40, and the MLS policy does not incur any degradation in the LER.
-Note that this estimation is optimistic since we assumed the number of rounds increases linearly with the undecoded sequence length – in reality, it could be worse since the decoder latency can increase exponentially with the undecoded sequence length.
-Figure 15. Reduction in the peak memory usage when using the Midpoint configuration with software offloading.
+Figure 14 shows the memory required for different workloads with the Midpoint configuration (memory requirements were determined using the code distances estimated by Azure QRE).
+Due to longer undecoded sequences, the MFD policy can require 3–4 orders of magnitude more memory than the MLS policy, which requires around 0.01MB of memory.
+The RR policy can require 3-10$\times$ more memory than the MLS policy.
+Figure 15. Reduction in the longest undecoded sequence length when using the Midpoint configuration with software offloading.
 […]
@@
 7.3.3. Software Offloading
-The longest undecoded sequences (and consequently the memory requirements) can be reduced further by leveraging software decoders.
-The only constraint while doing so is that due to longer software decoding latencies, critical decodes should not be delayed because prior software decodes for a logical qubit have not yet finished.
-For evaluating the effect of software decoding, we set the number of hardware decoders to the Midpoint configuration and make a pessimistic assumption that a single slice worth of syndromes takes three slices (about 3$\times$×d microseconds) to be decoded in software (in reality, it could be much lower with optimized software decoders (Higott2023, )).
-Fig.
-15 shows the reduction in the peak memory usage for all scheduling polices when software offloading is performed – software offloading can achieve a reduction of up to 3x.
+The longest undecoded sequences can be reduced further by leveraging software decoders.
+The only constraint while doing so is that critical decodes should not be delayed because prior software decodes for a logical qubit have not yet finished.
+For evaluating the effect of software decoding, we set the number of hardware decoders to the Midpoint configuration and make a pessimistic assumption that a single slice worth of syndromes takes three time slices (about 3$\times$×d microseconds) to be decoded in software (in reality, it could be much lower with optimized software decoders (Higgott and Gidney, 2023)).
+Figure 15 shows the reduction in the longest undecoded sequence length for all scheduling polices when software offloading is performed – software offloading can achieve a reduction of more than 30%.
+Figure 16.
+Total reduction in decoders required when considering 15-to-1 MSD factories with the Midpoint configuration, EDPC layout.
+The number of MSD factories was fixed at 24 for all workloads, and the number of factories estimated by Azure QRE are annotated below the benchmark names.
+[…]
+[some figure]
+7.3.4. Incorporating Magic State Distillation
+So far, our evaluations have considered only algorithmic logical qubits.
+Since Lattice Surgery Compiler takes the qubit layout as an input, we fixed the number and locations of magic state distillation (MSD) factories in the EDPC layout at 24 for all workloads.
+This number was chosen to maintain a uniform EDPC layout across all workloads, and it is also the median number of MSD factories estimated by Azure QRE (annotated below the benchmarks names in Figure 16).
+15-to-1 distillation is mapped to a Compact layout (Litinski, 2019a).
+Using this layout and the QASM circuit for a 15-to-1 MSD factory, the MLS policy was used to virtualize decoders.
+Table 1 shows the reduction in the number of decoders and the longest period for which a qubit remains undecoded for a single distillation factory.
+A 15-to-1 distillation factory requires five logical qubits (Litinski, 2019a) and with two decoders each for $X/Z$ observables if all qubits are supplied their own decoders.
+This can be reduced by 60% to 4 decoders with virtualization.
+Table 1. Virtualizing decoders using the MLS policy can reduce the decoders required for 15-to-1 distillation without a substantial backlog of undecoded slices.
+All Qubits After Virtualization
+Longest Undecoded
+Sequence Length
+15-to-1 Distillation 10 decoders 4 decoders 4 time slices
+Assuming that decoders are independent for every MSD factory, the reduction in the number of decoders can be computed using data from Figure 12 and Table 1, as shown in Figure 16.
+The MLS policy can reduce the total number of decoders required by up to 4$\times$.
+MSD is a probabilistic process, and factories will not be in sync.
+at all times.
+Thus, treating decoders for all factories to be independent is a conservative setting – in reality, it is likely that decoders can be virtualized among multiple factories, which can yield further reductions in the number of decoders in the system.
+7.3.5. Dynamic Scheduling
+Periodic events such as heating effects, cosmic rays, can result in a temporary increase in the error rate for some qubits.
+Such error bursts can necessitate dynamic scheduling of decoding for affected qubits.
+For evaluating the overheads of such error bursts on the virtualized decoders, we configured a burst probability $P_{Burst}$.
+For a workload requiring $N$ slices, $P_{Burst}\times N$ slices are randomly selected from which an active logical qubit is selected with a probability of $P_{Burst}$ to be decoded.
+Figure 17 shows the increase in the decoders required for different $P_{Burst}$ – dynamic scheduling only increases the decoder resource requirements for a 10% probability of a burst error happening.
+This shows that virtualization is effective even with moderate error burst probabilities.
+Figure 17. Normalized increase in the number of decoders required with the Midpoint configuration due to error bursts with probability $P_{Burst}$.
+[…]
+[some figure]
+Figure 18.
+(a) Total decoding task required to process the longest undecoded sequence of syndromes for all workloads, with the normalized decoder latency for the Surface Code (SC) as 0.5 and 0.99 for the qLDPC code.
+The Midpoint configuration was used for both cases; (b) Relative increase in the LER for different workloads for the SC and qLDPC setups.
+[…]
+[some figure]
+7.3.6. Slowdown due to Decoding Latencies
+Do slow and complex decoders, such as the ones used for qLDPC codes, work well with virtualization?
+Processing undecoded syndromes requires an additional number of rounds that is dependent on the decoding processing rate.
+For the Surface Code (SC), decoders that process syndromes significantly faster than the syndrome generation rate have been proposed (Vittal et al., 2023; Barber et al., 2023; Ziad et al., 2024).
+However, unlike the Surface Code, quantum LDPC codes (Bravyi et al., 2024; Roffe et al., 2020; Gong et al., 2024; Hillmann et al., 2024; Wolanski and Barber, 2024) have considerably slower decoders with software latencies in the order of milliseconds.
+To put that in context, software decoders for the Surface Code have microsecond latencies (Higgott and Gidney, 2023).
+Assuming accelerated decoders for qLDPC codes are designed, it is reasonable to expect them to be slower than Surface Code decoders.
+To see the impact of such slow decoders on virtualization, we modified our framework to support heterogeneous systems consisting of qLDPC blocks and Surface Code patches for compute (Stein et al., 2024).
+To do this, we assumed every magic state consumed involved an additional decoder to decode the ancillary system required for computation with qLDPC codes (Cross et al., 2024).
+[…]
+The results above show that the benefits of virtualization are only apparent when the decoding latency is much smaller than the syndrome generation latency.
+For both qLDPC and Surface Codes, it is thus highly desirable to have short decoding latencies, since this can enable more efficient virtualization.
+Decoders thus need not be designed for resource efficiency – they can be built for speed and accuracy.
+Resource savings can be attained at a system-level via virtualization.
 7.4. Discussion
-The results shown in previous sections show that VQD reduces the number of hardware decoders for algorithmic logical qubits by nearly one order of magnitude for most workloads with the Midpoint configuration, which, when combined with the MLS scheduling policy, results in significantly reduced memory requirements and low undecoded sequence lengths.
-Memory requirements and undecoded sequence lengths can be further reduced by offloading some non-critical decodes to software.
-[…]
-Applicability to other codes: While our evaluations have focused on the Surface Code, other error-correcting codes such as quantum LDPC codes (Bravyi2024, ) also require decoding with a significantly higher complexity than Surface Code.
-qLDPC codes use Belief Propagation for decoding errors (Old2023, ) along with Ordered Statistics Decoding (OSD) (Roffe2020, ; Fuentes2021, ) to enable accurate decoding.
-These algorithms are complex and highly resource intensive (Gong2024, ).
-Building fast, accurate decoders for such codes will likely require significant hardware resources.
-Our work enables amortizing the cost of expensive decoders via virtualization to build efficient and scalable quantum memory using qLDPC codes.
-Better Capacity Planning: We envision large-scale quantum computers will be closely integrated with HPC-style systems, where scientific applications can leverage quantum subroutines using QPUs (humbleQPU, ).
+The results shown in previous sections show that VQD reduces the number of hardware decoders for algorithmic logical qubits by nearly one order of magnitude for most workloads with the Midpoint configuration, which, when combined with the MLS scheduling policy, results in significantly small undecoded sequence lengths of $\sim$10 time slices.
+With the inclusion of distillation factories, the total reduction in the number of decoders is up to 4$\times$.
+Faster or Resource Efficient Decoders?: The faster a decoder is in relation to the syndrome generation latency, the more effective is decoder virtualization.
+Decoders with smaller hardware footprints (Vittal et al., 2023; Barber et al., 2023) generally sacrifice decoding accuracy for speed and hardware-efficiency.
+This work indicates that hardware-efficiency can be achieved with effective decoder virtualization on a system-level rather than at a module-level by leveraging fast, accurate decoders that need not be hardware efficient by themselves.
+This lends more motivation to the design of neural network decoders (Bausch et al., 2023; Acharya et al., 2024), which can be resource intensive but offer code distance independent decoding latencies and can leverage existing hardware (CPUs/GPUs) rather than require custom FPGA/ASIC solutions.
+This insight can be applied to build fast and accurate decoders for other codes such as qLDPC, Color codes as well.
+Better Capacity Planning: We envision large-scale quantum computers will be closely integrated with HPC-style systems, where scientific applications can leverage quantum subroutines using QPUs (Britt and Humble, 2017).
 In this setting, non-critical software decoders can run on traditional HPC platforms to alleviate the pressure on hardware decoders.
@@
 Prior to this work, Bombín et al.
-(Bombin2023, ) introduced modular decoding, which is the closest work that divides the global decoding task to sub-tasks without sacrificing decoder accuracy.
-However, this work, and other works such as parallelized window decoding (Skoric2023, ; Tan2022, ) always assume decoders for every logical qubit.
+(Bombín et al., 2023) introduced modular decoding, which is the closest work that divides the global decoding task to sub-tasks without sacrificing decoder accuracy.
+However, this work, and other works such as parallelized window decoding (Skoric et al., 2023; Tan et al., 2022) always assume decoders for every logical qubit.
 This work shows that not all qubits require access to fast decoders at all times, thus allowing decoders to be virtualized.
@@
 System-level Studies: Delfosse et al.
-(Delfosse2023, ) studied the speed vs.
+(Delfosse et al., 2023) studied the speed vs.
 accuracy tradeoff for decoders used in FTQC.
-XQSim (Byun2022, ) is a full-system FTQC simulator.
+XQSim (Byun et al., 2022) is a full-system FTQC simulator.
 Stein et al.
-(Stein2023, ) proposed a heterogeneous architecture for FTQC, virtual logical qubits were proposed in (Baker2021, ), Lin et al.
-(Lin2023, ) explored modular architectures for error-correcting codes and scheduling for distillation factories was proposed in (Ding2018, ).
-(Kim2024, ) described a blueprint of a fault-tolerant quantum computer.
-Decoder Designs: Neural network based decoders (googleRNN, ; ueno2022, ; Overwater2022, ; Meinerz2022, ; Gicev2023, ; Varsamopoulos2017, ), LUT-based decoders (Das2022lilliput, ; Tomita2014, ), decoders based on the union-find algorithm (Das2022afs, ; Riverlane2023, ), and optimized MWPM decoders (Vittal2023, ; Alavisamani2024, ) have been proposed.
+(Stein et al., 2023) proposed a heterogeneous architecture for FTQC, virtual logical qubits were proposed in (Baker et al., 2021), Lin et al.
+(Lin et al., 2023) explored modular architectures for error-correcting codes and scheduling for distillation factories was proposed in (Ding et al., 2018).
+(Kim et al., 2024) described a blueprint of a fault-tolerant quantum computer.
+Decoder Designs: Neural network based decoders (Bausch et al., 2023; Ueno et al., 2022; Overwater et al., 2022; Meinerz et al., 2022; Gicev et al., 2023; Varsamopoulos et al., 2017; Acharya et al., 2024), LUT-based decoders (Das et al., 2022a; Tomita and Svore, 2014), decoders based on the union-find algorithm (Das et al., 2022b; Barber et al., 2023; Ziad et al., 2024), and optimized MWPM decoders (Vittal et al., 2023; Alavisamani et al., 2024) have been proposed.
 In general, neural network decoders are generally far slower and therefore not ideal for fast qubit technologies such as superconducting qubits.
-Other predecoders (Delfosse2020, ; Smith2023, ) and partial decoders (Caune2023, ) have also been proposed.
-Decoders based on superconducting logic (Ueno2021, ; Ravi2023, ) target cryogenic implementations.
+Other predecoders (Delfosse, 2020; Smith et al., 2023) and partial decoders (Caune et al., 2023) have also been proposed.
+Decoders based on superconducting logic (Ueno et al., 2021; Ravi et al., 2023) target cryogenic implementations.
 9. Conclusions
@@
 In this work, we take a full-system view and show that with the right decoder scheduling policy, it is not necessary for an error-corrected quantum computer to provide every logical qubit with a dedicated hardware decoder.
-The MLS policy enables the reduction of hardware decoders by up to 10x while requiring $\sim$100 MB or less of memory for storing undecoded syndromes without increasing the program execution time or the target logical error rate.
+The MLS policy enables the reduction of hardware decoders by up to 10x without increasing the program execution time or the target logical error rate.
 The efficacy of the MLS policy is enhanced with software offloading of some decoding tasks.
 We also propose a noise-adaptive scheduling mechanism that can prioritize the decoding of logical qubits that incur a temporary increase in the physical error rate.
+Via virtualization, decoder systems can be made more scalable, and this insight can allow algorithm and hardware designers to focus on making decoders faster and more accurate, no matter the hardware cost.